### v1

# Deep Learning for Pathological Speech: A Survey

Abstract
Advancements in spoken language technologies for neurodegenerative speech disorders are crucial for meeting both clinical and technological needs. This overview paper is vital for advancing the field, as it presents a comprehensive review of state-of-the-art methods in pathological speech detection, automatic speech recognition, pathological speech intelligibility enhancement, intelligibility and severity assessment, and data augmentation approaches for pathological speech. It also highlights key challenges, such as ensuring robustness, privacy, and interpretability. The paper concludes by exploring promising future directions, including the adoption of multimodal approaches and the integration of graph neural networks and large language models to further advance speech technology for neurodegenerative speech disorders.
Index Terms:
Pathological speech, neurodegenerative speech disorders, speech processing, deep learning.
I Introduction
Speech production is a complex mechanism involving cognitive planning, coordinated muscle activity, and sound creation [1]. The process starts in the brain with the conceptualization of a message, followed by the organization of phonetic and prosodic plans such as rhythm and style. The motor cortex then orchestrates the activation of approximately $100100$ muscles, enabling the articulatory organs such as the tongue, lips, and jaw, to shape the vocal tract and produce specific sounds. The initial sound is generated in the larynx, where air from the lungs causes the vocal folds to vibrate. These phonatory structures adjust voice quality and prosody, while the articulatory organs further refine the sound by altering the shape of the vocal tract. Finally, the resulting speech is emitted through the oral and nasal cavities. Given the intricate coordination required for this process, any disruption to these finely tuned mechanisms can severely impair communication.
Neurodegenerative disorders such as Parksinson’s disease (PD), Amyotrophic Lateral Sclerosis (ALS), or Alzheimer’s disease are leading causes of such disruptions [2, 3, 4]. These disorders impair the brain regions and motor systems responsible for initiating, planning, and controlling the movements needed for speech production, resulting in a variety of speech disorders such as dysarthria, aphasia, apraxia of speech, dysphonia, or stuttering [5, 6, 7, 4, 8, 9]. Dysarthria and apraxia of speech, commonly seen in PD and ALS, are characterized by articulation deficiencies, vowel distortions, reduced loudness variation, hypernasality, or syllabification [5, 4]. Dysphonia, also frequent in PD and ALS, is marked by abnormal voice quality such as hoarseness and breathiness [6]. In contrast, aphasia typically presents as difficulties with word-finding, sentence construction, and understanding spoken or written language, and is most commonly associated with Alzheimer’s disease or other forms of dementia [8]. Lastly, stuttering, characterized by disruptions in the flow of speech such as repetitions or prolongations can also arise in neurodegenerative disorders like Parkinson’s disease [9].
Because of increasing population numbers and aging, the prevalence of neurological disorders, and hence, of these various speech disorders, is also rapidly increasing. In $20192019$, the World Health Organization estimated that over $8.58.5$ million people worldwide were living with PD [10], up from $6.16.1$ million in $20162016$ [11] and $2.52.5$ million in $19901990$ [11]. By $20402040$, this figure is projected to surpass $1717$ million [11]. Similarly, dementia affected more than $4646$ million people globally in $20152015$ and this figure is expected to rise to $131.5131.5$ million by $20502050$ [12]. The prevalence of ALS is also growing significantly, with cases anticipated to increase by nearly 70% between $20152015$ and $20402040$ [13]. This increasing prevalence of neurological disorders, and consequently of the associated speech disorders, underscores the need to prioritize speech disorders both in the context of clinical practice as well as in the context of speech-based technologies.
In clinical practice, diagnosing the presence of a speech disorder (i.e., discriminating between neurotypical and impaired speech) is crucial, since the presence of a disorder can be one of the earliest signs of neurodegenerative disorders [14, 15, 16]. Further, an accurate differential diagnosis of the speech disorder (e.g., discriminating between dysarthria and apraxia of speech) is also important, since it can provide clues about the underlying neuropathology [17, 18]. Additionally, monitoring speech characteristics such as severitys and intelligibility after diagnosis is essential for tracking disease progression and assessing the effectiveness of speech therapy interventions over time [2, 19]. Traditionally, clinicians conduct such evaluations through costly and time-consuming auditory-perceptual assessments as depicted in Fig. 2. However, because of the difficulty of detecting clinical-perceptual characteristics by ear (particularly in the presence of mild impairments) and because the clinical-perceptual characteristics of several speech disorders (e.g., of dysarthria and apraxia of speech) overlap, diagnosing the presence of a disorder and discriminating between various disorders is hard for non-expert clinicians and even expert inter-rater agreement can be low [20, 21]. In the context of speech-based technologies, patients with speech disorders often struggle to access and effectively use various virtual assistant tools like Cortana, Alexa, and Siri [22]. As speech disorders become more prevalent with neurodegenerative conditions, it is critical to prioritize both the diagnosis and treatment of these disorders in clinical practice, while also ensuring that patients with such impairments have equitable access to speech-based technologies. Addressing these barriers could lessen the burden on the health care system and significantly improve the patients’ quality of life and their ability to engage with everyday digital tools.
Figure 1: Traditional auditory-perceptual assessment in clinical practice. The clinician listens to the (potential) patient speaker and assesses by ear the various characteristics of the speech.
Figure 2: Automatic pathological speech analysis system. The automatic model is trained to detect and analyze speech impairments. Clinicians may use the insights provided by the automatic model to organize therapeutic sessions accordingly. Additionally, they may perform manual acoustic analysis to gain further insights into the patient’s speech patterns and impairments, providing complementary information to the automatic model’s decision.
Aiming to assist the clinical diagnosis and treatment of patients suffering from neurodegenerative disorders, there has been a growing interest in the research community to develop automated and objective methods for pathological speech analysis as shown in Figure 2. These advanced technologies are designed to minimize bias, enhance diagnostic accuracy, and streamline the assessment process, ensuring greater efficiency and consistency. Clinicians can then use insights provided by the automatic models to organize therapeutic sessions accordingly, ensuring that the treatment addresses the specific impairments identified. Further, clinicians may perform additional manual acoustic analysis to gain further insights into the patient’s speech patterns and impairments, providing complementary information to the automatic model’s decision for a comprehensive therapeutic approach. Besides the clinical domain, efforts have been directed towards developing various speech-based technological applications aimed at pathological speakers, such as automatic speech recognition systems (ASRs) [23], speech synthesis systems [24, 25, 26, 27] or intelligibility enhancement solutions [28, 29, 30, 31]. To the best of our knowledge, there is currently no comprehensive survey paper discussing the research directions and challenges of this area both from a clinical as well as a technological perspective. Although there have been related works such as [32], they primarily provide an overview of acoustic-articulatory features in neurodegenerative disorders. Other studies such as [33, 34, 35, 36, 37] mainly focus on the discrimination between neurotypical and impaired speech. In $20162016$, [38] discussed some of the wider challenges faced in the pathological speech domain. However, this work addresses only a limited set of challenges and is now somewhat outdated, given the rapid advancements in the field over the past decade. In contrast, as depicted in Figure 3, our work aims to fill this gap by providing an extensive review of the field encompassing pathological speech from various clinical and technological perspectives, such as automatic discrimination between neurotypical speech and speech disorders, ASR systems for pathological target speakers, enhancement systems aiming to enhance the intelligibility of pathological speech, severity and intelligibility assessment and data augmentation approaches.
Figure 3: Overview of Deep Learning for Pathological Speech: This framework leverages various input features and datasets processed through deep learning models for tasks including detection, recognition, enhancement, assessment and explainaibility of pathological speech. It employs a range of metrics to evaluate model performance, providing a comprehensive approach to improving speech impairment detection and treatment.
The remainder of the paper is organized as follows. Section II provides a high level overview of the handcrafted acoustic features used in pathological speech analysis. Section III describes various accessible and non accessible pathological datasets employed in the literature. Section IV describes the different approaches to automatic pathological speech detection. Section V examines pathological speech in the context of ASR systems. Section VI discusses speech enhancement techniques aimed at improving pathological speech intelligibility. Section VII describes approaches for automatically estimating the severity and intelligibility of pathological speakers. Section VIII summarizes the various data augmentation methods used in automatic pathological speech systems. The robustness issues arising in this field are discussed in Section IX-A. Finally, Section X explores the interpretability of pathological speech models. In summary, our contributions are the following:
  1. 1.
    We present the first comprehensive survey of automatic approaches for pathological speech from a clinical and technological perspectives ranging from detection, recognition, enhancement, assessment and interpretability.
  2. 2.
    We highlight current limitations in the field and propose several promising future directions for research in pathological speech processing.
II Pathological Speech: Acoustic Characteristics and Representation
Pathological speech exhibits a range of acoustic anomalies, including deviations in pitch, loudness, vowel space reduction, and articulation [39, 40]. Additionally, it can lead to asymmetrical tension in the vocal folds, resulting in irregular vibrations and, consequently, an abnormal fundamental frequency. Excessively high or low fundamental frequency, combined with excessive vocal intensity, can exacerbate the severity of pathological voice conditions, producing characteristics such as shrillness, screechiness, hoarseness, or huskiness [41]. Moreover, patients (such as those diagnosed with PD) can also show inconsistent rhythmic structures in comparison to control groups under various auditory feedback conditions [42]
To capture such abnormalities as biomarkers for automatic pathological speech detection, researchers have employed various handcrafted acoustic features such as OpenSMILE [43], Mel-frequency cepstral coefficients (MFCCs) [44], or spectro-temporal sparsity features [45]. Further, various time-frequency representations such as the short-time Fourier transform (STFT) or the Mel spectrogram have also been directly exploited in combination with deep learning approaches to directly learn pathology-discriminant cues [46]. More recently, latent embeddings derived from self-supervised models have been used as more powerful representations of speech patterns and the various impairments [43, 47, 48]. It is important to note that a significantly larger set of features and representations have been employed to characterize disorders across various pathologies than the ones outlined above. However, here we focus on highlighting the most commonly utilized ones, which will be further elaborated upon in the following.
II-A Handcrafted Acoustic Features
OpenSMILE
Among the various features used in pathological speech detection, OpenSMILE features have been widely explored in the literature in conjunction with traditional machine learning algorithms [49, 50, 51, 52, 31, 53, 54, 55, 56, 57, 58, 59, 60, 61]. The OpenSMILE feature set includes a 6552-dimensional feature vector that primarily consists of low-level audio features such as CHROMA, CENS, loudness, MFCCs, and other spectral features. In the context of pathological speech, these features can capture the subtle anomalies in speech patterns that are indicative of disorders. OpenSMILE’s comprehensive feature set allows for the analysis of prosodic elements like pitch, jitter, shimmer, and formant frequencies, which are crucial for identifying and differentiating various speech pathologies. However, OpenSMILE features are very general features that have been used for a variety of speech applications such as emotion recognition [62], speech recognition [63] and they are not specifically handcrafted to capture pathological cues, thus may not be the most suitable features for pathological speech analysis.
MFCCs
MFCCs have garnered significant attention in pathological speech detection due to their potential to characterize articulation deficiencies [64, 65, 66, 67, 68, 69, 70, 71, 72]. While MFCCs have demonstrated promising results, they exhibit high sensitivity to background noise and present challenges in real-time applications due to their complex nature and lack of interpretability.
Spectro-temporal sparsity
Since pathological speech can be breathy, semi-whispery, and is characterized by abnormal pauses and imprecise articulation, it can be expected that its spectro-temporal sparsity differs from the spectro-temporal sparsity of neurotypical speech. To characterize spectro-temporal sparsity, various sparsity-based features have been introduced in [45, 73]. Although such features have been shown to be discriminative of various speech disorders[45, 73, 61, 74], they are highly sensitive to environmental artefacts such as noise and reverberation.
II-B Time-Frequency Representations
Input representations such as the STFT and its variants allow for the analysis of speech signals in both time and frequency domains, providing valuable insights into speech characteristics such as pitch, formant shifts, and spectral irregularities. The STFT is often employed due to its ability to capture dynamic changes in the signal over time, which is critical for identifying variations in speech patterns linked to speech disorders [75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 25, 85, 86]. Furthermore, wavelet transforms and the Continuous Wavelet Transform offer better time-frequency localization, which is especially beneficial for analyzing transient and non-stationary features of pathological speech [87].
II-C Self Supervised Embeddings
Even though handcrafted acoustic features and time-frequency input representations have achieved promising results in the analysis of pathological speech, their performance remains limited. The current state-of-the-art input representations for pathological speech primarily consist of latent embeddings derived from self-supervised models such as wav2vec 2.0 (w2v2) [88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 47, 101, 102, 103, 104, 61, 105, 106, 107, 108, 109, 110, 111]. These advanced models have proven to be more effective in capturing complex speech patterns and nuances, leading to improved performance in tasks like detection and recognition [112, 48], However, there is still ongoing research to further refine these embeddings and explore additional models to enhance their robustness and generalization across diverse pathological speech conditions [113, 114, 115].
III Pathological Speech Datasets
Datasets for pathological speech research remain scarce due to several inherent challenges in data collection. One of the primary difficulties is the sensitive nature of the population involved. Recruiting participants suffering from neurological disorders requires careful consideration of ethical and privacy concerns, as well as navigating the potential stigma associated with such disorders. Additionally, the physical and cognitive challenges faced by these individuals can complicate the process of obtaining high-quality speech recordings. Another key challenge is the variability in pathological speech patterns. Speech impairments can manifest in numerous ways and may vary widely across individuals, even within the same diagnostic category. This variability makes it difficult to create standardized protocols for data collection that ensure consistency and relevance across samples. Moreover, speech impairments may fluctuate over time, further complicating the process of capturing representative speech samples. When collecting datasets for pathological speech research, it is essential to prioritize inclusivity, ensuring a diverse representation of different speech disorders, age groups, and demographics. Additionally, data collection protocols should consider the comfort and cooperation of participants. Ethical considerations must also be at the forefront, ensuring informed consent and the protection of participant privacy. Lastly, it is important to design flexible and scalable collection methods that can capture a range of speech characteristics while maintaining consistent quality across diverse individuals and conditions. In the remainder of this section, we briefly review pathological datasets commonly used in the literature. Please note that due to restricted access to most of the pathological neurodegenrative datasets, we are unable to provide detailed information such as duration, number of male and female speakers, etc., for all of them. A summary of these datasets and their characteristics is presented in Table I.
TORGO  [116].  The TORGO dataset contains speech recordings from English speaking subjects (controls and patients) and the corresponding three dimensional articulatory features. The patients suffer from ALS or Cerebral Palsy (CP). The dataset consists of seven (three females, four males) controls and eight (three females, five males) dysarthric speakers respectively.
PC-GITA [117].  The PC-GITA dataset is a Spanish dataset containing speech recordings from $5050$ controls and $5050$ patients suffering from PD. The two groups of speakers are age- and gender-matched, with $2525$ female and $2525$ male speakers in each group.
MoSpeeDi [118].  The MoSpeeDi dataset is a French dataset containing speech recordings from $466466$ controls and $138138$ patients suffering from various types of motor speech disorders such as dysarthria or apraxia of speech. While subgroups of age- and gender-matched controls and patients can be found within the database, the overall dataset is not age- and gender-matched.
Nemours [119].  The Nemours dataset is an English dataset containing speech recordings from $1111$ male speakers with varying degrees of dysarthria severity. This dataset includes a total of 74 sentences and 814 short nonsense sentences.
CUDYS [120].  The Cantonese Dysarthric Speech Corpus (CUDYS) contains speech recordings from $55$ controls and $1111$ dysarthric patients. The control group consists of $22$ female and $33$ male speakers whereas the dysarthric group consists of $55$ female and $66$ male speakers. The dataset has a total duration of $1010$ hours, with the speakers’ ages ranging between $2525$ and $6767$ years.
AMSDC [121].  The Atlanta Motor Speech Disorders Corpus (AMSDC) is an English dataset containing speech recordings from $9999$ ($6262$ males and $3737$ females) patients diagnosed with neurogenic disorders. This dataset is essential for studying motor speech disorders, though access to the full dataset might be restricted.
EST [122].  The EST dataset is a Dutch dataset containing speech recordings from 16 dysarthric patients ($1616$ males). The dataset has a total duration of $\approx$ $66$ hours and the speakers’ ages range between $3434$ and $7575$ years.
EasyCall [123].  The EasyCall dataset is an Italian dataset containing speech recordings from $2424$ controls and $3131$ patients diagnosed with PD, Huntington’s disease, ALS, and peripheral neuropathy. The control group consists of $1010$ female and $1414$ male speakers whereas the patient group consists of $1111$ female and $2020$ male speakers. The corpus includes a total of 21,386 speech commands, with 10,077 recordings from healthy subjects and 11,309 from dysarthric patients.
COPAS [124].  The Corpus of Pathological and Normal Speech (COPAS) contains speech recordings from Dutch speakers with $197197$ pathological speakers and $122122$ normal speakers. The database comprised of 8 distinct pathological categories.
ParkCeleb [125].  The previously reviewed datasets are not longitudinal and do not allow tracking the progression of the speech disorder within the same patient along time. To address this gap, the English ParkCeleb dataset was recently introduced in [125]. This dataset contains speech recordings from $4040$ controls and $4040$ patients suffering from PD.
Summary of datasets.  As outlined in Table I, many of the pathological speech datasets used in the literature are private and not openly accessible, limiting their availability for broader research and replication. Furthermore, datasets such as TORGO, although publicly available, have been shown to contain detrimental recording artefacts [43], which limits their suitability for studying pathological speech characteristics or for developing automatic pathological speech detection approaches [43, 113]. This lack of accessible, high-quality data creates a substantial barrier for developing and validating effective algorithms and tools. Consequently, the advancement in understanding and treating pathological speech conditions is slowed, limiting the potential for innovative solutions and breakthroughs in this critical area of healthcare. Ensuring the availability of clean, reliable, and comprehensive datasets is essential to foster meaningful research and development. Additionally, many existing pathological speech datasets are predominantly in English, restricting their applicability to non-English-speaking populations and limiting cross-linguistic comparisons. This linguistic bias creates challenges when attempting to develop universal speech models or tools that can be applied globally. Furthermore, most of these datasets are relatively small, often comprising only a limited number of participants, which hinders the development of robust, generalizable models. The small size of these datasets also makes it difficult to account for the variability in speech disorders across different populations and conditions. Another important limitation is the imbalance in demographic representation in terms of gender. Many datasets suffer from a skewed representation, often over-representing one gender, which can lead to biases in speech models. Additionally, some datasets may not include a sufficient variety of age groups or pathological conditions, further restricting their generalizability. These factors collectively underscore the need for high quality, larger, more diverse, and well-balanced pathological speech datasets that reflect the complexity and diversity of real-world populations.
Figure 4: Automatic Pathological Speech Detection Pipeline: The process begins with recording the patient’s speech, which is then processed to extract features that are meaningful and suitable for an AI classifier. These features are transformed into a feature space, where a neural network model analyzes them to make a diagnostic decision. This pipeline streamlines the detection process, enabling efficient and accurate identification of pathological speech patterns.
IV Automatic Pathological Speech Detection
TABLE I: Overview of Pathological Speech Databases. (PD: Parkinson’s disease, CVA: Cerebrovascular accident, TBI: Traumatic brain injuries, ALS: Amyotrophic lateral sclerosis, Huntington’s disease, 3D: Articulatory (3D), L: Longitudinal data, Accessible: Accessible through a license for research purposes )
Database Number of Speakers Type of Impairment Language Modality Accessibility
Control Pathological
PC-GITA 50 50 PD Spanish Speech Accessible
TORGO 07 08 ALS/CP English Speech + (3D) Public
MoSpeeDi 466 138 Dysarthira/Apraxia French Speech Not Accessible
Nemours - 11 Dysarthira English Speech Not Accessible
CUDYS 05 11 Spino-cerebellar ataxia Cantonese Speech Not Accessible
AMSDC 62 37 CVA, PD English Speech Not Accessible
Dutch EST - 16 Dysarthria, TBI Dutch Speech Not Accessible
EasyCall 24 31 PD, HD, ALS Italian Speech Public
COPAS 197 122 Dysarthira and others Dutch Speech Accessible
ParkCeleb 40 40 PD English Speech (L) Accessible
Neurodegenerative diseases such as PD and ALS can significantly impact various sensorimotor and articulatory processes, leading to pathological speech impairments such as dysarthria, apraxia of speech, aphasia, dysphonia. As discussed in Section I, clinicians traditionally rely on laborious and time-consuming auditory-perceptual measures to diagnose these speech impairments accurately. To address the various challenges associated with auditory-perceptual assessments, there has been a growing interest in the research community to develop automatic approaches for diagnosing pathological speech. The goal of these approaches is to enhance the detection accuracy to match or surpass human accuracy, thereby ensuring more consistent and reliable diagnoses. Automated systems leverage advanced technologies such as classical machine learning and deep learning models with input representations as discussed in Section II to analyze speech patterns and identify anomalies indicative of these neurodegenerative conditions. In the following, we briefly summarize various classical machine learning-based and deep learning-based approaches. Additionally, insights on the speech modes used for pathological speech detection are provided.
IV-A Classical Machine Learning-based Approaches
[126, 127] employed random forests and support vector machines (SVM), reporting promising results on a small pool of speakers. Similarly, [128] demonstrated good performance in detecting pathological dysarthric speech in an Indian Tamil language dataset using Gaussian mixture model and hidden Markov model classifiers. [129] found that the use of articulatory features on a small set of 24 Czech speakers showed very good performance, but might not generalize well to larger sets of speakers. Despite the promising results from these methods using various spectral/cepstral features, it remains unclear which features are most discriminative in capturing pathological cues. [130] showed that the cepstral coefficients are more robust and achieved the highest performance on a Spanish dataset when used as input to the SVMs for detection of pathological speech. [131] exploits $63736373$ acoustic, $36003600$ articulatory, and $44$ sensory features for the detection of ALS using deep neural networks and showed state-of-the-art detection performance on a dataset consisting of $1111$ patients and $1111$ healthy speakers. In a similar fashion, [132] employed SVM with openSMILE features on a dataset of 123 Hebrew speakers. Additionally, [133] utilized glottal features and demonstrated that combining glottal features with openSMILE features enhances the performance of dysarthric speech detection. [45] introduced spectro-temporal sparsity features combined with SVMs, showing these features to be more effective discriminators for dysarthric speech compared to temporal sparsity features. [134] applied Grassmann discriminant analysis on spectro-temporal subspaces, using singular value decomposition based on clinical evidence of spectro-temporal distortions. Their models were evaluated on datasets-PC-GITA and MoSpeeDi.
The major limitation of such works is their focus on small speaker sets and bias towards factors such as gender, recording conditions, age, class, or language, leading to poor generalization in clinical scenarios. To investigate the impact of the distribution shift, [135] employed cross-database models for dysarthric pathological speech detection. They demonstrated that an SVM model trained on the UA-Speech dataset showed a significant performance drop when tested on the AMSDC dataset. This highlights the need for future work to develop models robust enough to handle distribution shifts, including variations in recording conditions and other factors. Furthermore, classical machine learning-based methods primarily rely on hand-crafted features, which may not effectively capture all abstract pathological cues. Finally, these methods also tend to overlook valuable metadata such as speaker identity, gender, and language.
IV-B Deep Learning-based Approaches
[136, 30, 137] utilized convolutional neural networks (CNN) for dysarthric speech detection on a dataset consisting of 26 speakers, including 13 patients diagnosed with ALS and their gender-matched counterparts respectively. Although this method demonstrated promising results, there is significant variation in pathological detection at the individual speaker level. [46] employed phase spectra such as modified group delay and instantaneous frequency spectra, demonstrating that using these phase spectra as complementary features to magnitude spectra improves the performance of CNN models in detecting dysarthric speech. Similarly, [56] employed temporal envelope features with CNN for dysarthric classification. [138] used SVM and deep neural networks for the detection of pathological speech on a dataset of speakers diagnosed with ALS. While most proposed approaches are single-task approaches aiming at achieving only pathological speech detection, [139] introduced the first ever multi-task system where they trained jointly pathological classification (primary task) and other 11 articulatory deficit tasks. The representations learned during joint training are constrained by various speech tasks and shows better generalization in pathological speech detection across speakers on a Spanish dataset.
Most studies to date have not exploited the temporal nature of audio frames. [140, 51, 141] addressed this issue by introducing a long short-term memory (LSTM) temporal network for dysarthric speech classification. [140] evaluated LSTM on a Mandarin dataset comprising of 69 speakers, 31 diagnosed with dysarthria and the remaining as normal speakers, while [51] evaluated an LSTM-based model on the TORGO dataset. Although the methodology in [140] showed promising results, the model has been evaluated only on the syllables making it harder to generalize across spontaneous speech.
Despite the promising results achieved by the proposed methods using handcrafted acoustic features and time-frequency representations, their detection capabilities still fall short of human-level performance and remain inadequate for clinical applications. Instead of handcrafted acoustic features or time-frequency representations, researchers endeavored to leverage raw speech data and applied self-supervised learning models like w2v2 to derive meaningful contextual representations, with the goal of improving the detection performance of pathological speech [60, 142, 43]. To the best of our knowledge, such models have become the state-of-the-art models in pathological speech detection, outperforming other automatic approaches based on handcrafted acoustic features or time-frequency input representations. However, although advantageous, these models currently lack interpretability insights that might be useful for a clinician. Moreover, the models developed thus far are largely language-dependent, as the linguistic context plays a critical role in pathological speech. Developing language-agnostic pathological detection methods would be highly beneficial, allowing for universal application across patients worldwide [143].
IV-C Impact of Speech Mode on Pathological Speech Detection
The large majority of the previously reviewed automatic pathological speech detection approaches have been proposed and validated on controlled speech tasks.
Controlled speech, also known as non-spontaneous speech, involves utterances produced within a structured context. In this mode, neurotypical speakers and patients are given a phonetically balanced, carefully crafted text and asked to repeat it. The goal is to elicit pathological biomarkers and cues through these controlled utterances, which can offer insights into motor planning and execution. Tasks in this mode may include reading from a text and repeating specific words.
In contrast to controlled speech, non-controlled or spontaneous speech consists of unplanned utterances that mirror real-world communication. This type of speech requires greater cognitive effort for planning, motor execution, and articulation, making it more likely to reveal natural pathological biomarkers. Examples include storytelling, recounting incidents, or engaging in everyday conversations. The reliance on controlled speech, with its structured and phonetically balanced content, may not always capture the full spectrum of pathological cues. Incorporating spontaneous speech could provide a more comprehensive understanding of pathological speech patterns. It seems plausible that developing pathological speech detection methods for spontaneous speech could be highly beneficial, given its ease of collection [61]. Research in this area could make such methods more practical and effective, as spontaneous speech naturally reflects real-world communication and may offer a richer set of pathological cues.
V Pathological Speech Recognition
Automatic speech recognition (ASR) systems have significantly advanced over the years, demonstrating impressive results in converting raw speech signals into their corresponding textual form. This has resulted in the widespread use of various interactive devices, including smartphones and voice assistants. However, they often fail to recognize low resource pathological speech, including speech from patients suffering from various neurodegenerative conditions such as PD or ALS [22].
To alleviate this problem, several attempts have been made in improving the performance of ASR systems for low resource pathological speech. For example, [144] proposed a two-step speaker adaptation method. Initially, a model trained on extensive neurotypical speech data is fine-tuned for dysarthric ASR. Subsequently, this fine-tuned model undergoes further adaptation to the specific dysarthric speaker. Additionally, [145] proposed the additive angular margin loss to address intra-class variation among dysarthric speakers, demonstrating promising results with Japanese speakers.  [146] also showed that personalized ASR systems fine-tuned on pathological speech exhibit better recognition performance compared to speaker-independent ASR systems. However, the reliance on specific speakers in personalized ASR systems poses challenges for generalization settings.  [147] utilize lattice-free maximum mutual information to mitigate insertion errors, which are otherwise prevalent due to the slow speaking rates of individuals with dysarthria. Due to the limited amount of available pathological speech data, the progress of ASR for pathological speech remains restricted. In a different approach, [148] utilized transfer learning and found that speaker-based data selection leads to negative transfer. They recommended using utterance-based data selection with an entropy distribution to enhance recognition.  [149] further implemented a multi-task learning (MTL) framework through auto-encoder joint learning, utilizing bottleneck features on out-of-domain data. Employing MTL in pathological ASR showed a lower word error rate compared to its single-task counterpart.  [150] demonstrated that state-of-the-art ASR systems for pathological speech are significantly impacted by phoneme inaccuracies. To address this, they proposed Speech Vision—a transfer learning paradigm that converts word utterances into visual feature representations, aiming to recognize the shape of the word rather than relying on phonemes. A comprehensive overview of ASR systems for pathological speakers in terms of progress and challenges is provided in [23].
Due to the large number of parameters in state-of-the-art ASR models, fine-tuning them on low-resource pathological speech data is highly expensive. However, leveraging fine-tuning techniques such as LoRA (Low-Rank Adaptation) [151, 152] offers a more efficient alternative, reducing computational costs and making training more feasible. Additionally, LoRA’s modular approach might possibly make it easier to adapt the fine-tuned models to new speakers, providing flexibility in real-world applications or clinical scenarios.
VI Intelligibility Enhancement
of Pathological Speech
In pathological speech, enhancement refers to the process of improving the quality and intelligibility of speech samples affected by various impairments. The benefits of pathological speech enhancement are manifold. Firstly, the improved quality of pathological speech can be integrated into ASR systems, potentially enhancing the recognition accuracy of such speech, thereby increasing the usability of these systems for individuals with speech impairments. Secondly, these approaches can serve as data augmentation methods (r.f. VIII) for multiple applications such as pathological speech detection and automatic speech recognition.
Progress in pathological speech enhancement is limited, however, several attempts have been made by various researchers as described in the following.  [24] restructured the trajectory of formants to more closely align with the desired sound targets, thereby enhancing the accuracy of the acoustic representation. This refined formant trajectory was subsequently utilized to generate more intelligible speech, significantly improving the clarity and quality of the unintelligible pathological speech.  [153] utilized a technique where the phase of the source signal was varied across six microphone arrays to synthesize pathological speech utterances. This approach effectively increased the data samples, which were later used for data augmentation in dysarthric ASR. [25] implemented a pause insertion model along with a severity level coefficient in a neural multi-task text-to-speech system, effectively replicating and synthesizing dysarthric speech. The synthesized speech was then utilized as data augmentation for improving dysarthric speech recognition.  [154] enhanced the intelligibility of dysarthric speech by applying a methodology that includes pronunciation correction, insertion of missing sounds, devoicing of unvoiced phonemes, tempo and spectral adjustment, and removal of unwanted disfluencies. [155] demonstrated that modifying the short-term spectral content on a frame-by-frame basis can significantly improve the word-level intelligibility for patients with dysarthria. However, despite the substantial intelligibility improvements, the method’s speaker dependency poses challenges for generalizing across different speakers. In a similar speaker-dependent study, [156] employed a Kepstrum-based approach. [157] utilized multiple pronunciation patterns along with genetic algorithms and hidden Markov models [26, 27] to enhance dysarthric speech.  [158] implemented a two-stage framework in which dysarthric speech was initially recognized using an hidden Markov-based network. Subsequently, a synthesis system adapted to the dysarthric speaker was employed to produce a new speech output by correcting the mispronounced words, thereby enhancing the intelligibility of the dysarthric speech.  [159] used dynamic time warping with CNNs to improve the speech intelligibility.  [160, 161] utilized linear predictive coding mapping in conjunction with dynamic time warping between dysarthric speech and neurotypical speech to enhance dysarthric speech intelligibility in an auditory feedback mechanism. [162] employed feature space mapping through constrained maximum likelihood linear regression, aligning acoustic features from dysarthric speakers to those of neurotypical speakers. This approach demonstrated improved intelligibility of dysarthric speech, as evidenced by the perceptual evaluation of speech quality test. A similar study was conducted by [76], where dysarthric speech features were enhanced using time-delay neural network denoising autoencoders in a speaker-independent fashion. However, only a small subset of speakers were used in the evaluation, which may not generalize well to a larger pool of speakers. Motivated by the human auditory perception system, [163] introduced perceptually enhanced single frequency cepstral coefficients to improve the intelligibility of speech in dysarthric patients using i-vectors and a Gaussian mixture model with a universal background model.  [30] examined the spectro-temporal representation and showed it can be useful in improving the intelligibility of speech. [31] argued that prior methods for speech intelligibility enhancement primarily focus on speaker-dependent systems, and they introduced speech posterior feature sets designed to capture dysarthric attributes while ignoring speaker-specific characteristics. However, this assertion is not entirely accurate, as recent research has demonstrated that dysarthric speech is closely intertwined with speaker attributes [75].
In a different approach, [164] integrated Transformer-TTS, CycleVAE-VC, and LPCNet to synthesize intelligible and robust dysarthric speech, although the resulting speech lacked naturalness. [28] formulated pathological speech enhancement as a style transfer problem and proposed a generative adversarial network (GAN)-based method for dysarthric speech enhancement. Similarly, [29] used CycleGAN for converting dysarthric speech to a more intelligible speech. [165] introduced DiscoGAN to improve the speech intelligibility of dysarthric patients. This method converts pathological speech to neurotypical speech through feature mapping learned during DiscoGAN training with an adversarial and construction losses [165]. The approach showed promising results, though further validation with a larger and more diverse patient group is necessary to ensure generalizability across various pathological conditions. [166] explored a related approach and found that time-stretching techniques for dysarthric speech outperform existing state-of-the-art generative models. Their study concluded that using time-stretching with MaskCycleGAN achieves significantly better performance in dysarthric speech enhancement. However, they noted that despite these improvements, the method is not yet suitable for practical applications.
Dysarthric speech datasets typically have a limited vocabulary, but using a conversion model can expand the set of out-of-vocabulary words, which can then be used for data augmentation.  [167] presented the first end-to-end voice conversion system that trains dysarthric speech along with its corresponding textual form using a knowledge distillation approach. This method enables the replication of speech from dysarthric speakers with improved pronunciation and more precise articulation. To expand out-of-vocabulary word set, [78] utilized a voice conversion model to transform natural speech into synthetic dysarthric speech utterances. A detailed study of pathological voice conversion can be found in [168].
Neural encoder-decoder models have been shown to significantly outperform GAN-based systems in enhancing the intelligibility of dysarthric speech. However, their reliance on multi-step modules often results in inefficient and suboptimal perceptual quality. To address this issue, [169] proposed Unit-DSR, a self-supervised learning approach for dysarthric speech synthesis. It has been shown that the adversarial training framework is highly sensitive to training parameters, often leading to instability and model divergence[170, 171]. In contrast, diffusion models [172] show significant promise for applications in pathological speech enhancement, synthesis, and voice conversion. On a similar lines,  [173] employed the Glow-TTS diffusion model for generating dysarthric speech and used that as data augmentation for training a dysarthric ASR system.
VII Intelligibility and Severity Assessment
of Pathological Speech
Intelligibility.  Intelligibility of pathological speech is a critical indicator for evaluating the effectiveness of speech therapy and tracking the progression of various disorders. To reduce the burden of evaluating pathological speech intelligibility in clinical practice, automatic approaches have been proposed in the literature. Automatic pathological speech intelligibility assessment methods are typically categorized into two main approaches, i.e., blind and non-blind approaches [75]. In blind approaches, the objective is to assess the intelligibility of impaired speech without exploiting reference neurotypical speech data [174, 175, 176, 177, 178, 179, 180]. These approaches primarily focus on extracting acoustic features such as jitter, fundamental frequency, shimmer, formant frequencies, etc., that are believed to be closely correlated with speech intelligibility. These features are then used in regression models to estimate the intelligibility of pathological speech. Non-blind approaches, by contrast, rely on intelligible speech from neurotypical speakers as a basis for estimating the intelligibility of pathological speech [181, 182, 183, 184, 185, 186, 187, 188, 189, 190]. Such approaches typically use features extracted from ASR systems - which have been trained on large amounts of neurotypical speech - to train regression models to estimate the intelligibility of pathological speech. To avoid the burden of collecting and transcribing a large amount of neurotypical speech data required for such systems, [191] proposed the pathological short-time objective intelligibility measure (P-STOI) adapted from the speech enhancement domain. The P-STOI measure first calculates an utterance-dependent fully intelligible representation from a small set of neurotypical speakers. The intelligibility of the pathological utterance is then evaluated by quantifying its divergence from this reference representation in terms of the short-time spectral correlation. While advantageous, P-STOI requires recordings of the same utterance from intelligible neurotypical speakers, which may not always be available. To mitigate this issue, [192] developed a method to generate synthetic reference speech for assessing pathological speech intelligibility. Likewise, [193] introduced a method for invariant reference speaker pair selection to evaluate speech intelligibility. This approach leverages the SpeechSplit framework, using dynamic time warping to measure the divergence between latent representations of healthy and pathological speech utterances. In a different approach, [194] introduced sub-space based intelligibility measures based on the premise that dominant spectral patterns in pathological speech deviate significantly from those of intelligible speech. Although such measures result in a lower performance than measures exploiting neurotypical intelligible speech, they can be directly used in practical scenarios where such speech material in not available or easy to generate.
Severity.  Besides intelligibility assessment, severity assessment is another important research area where developing tools for this purpose could greatly assist in automatizing the tedious process of screening patients and categorizing them into different subgroups based on the severity level. The methods developed in the literature for this purpose can be categorized into two categories, i.e., traditional machine learning-based approaches using the Mahalanobis distance classifier [174], SVMs [195], GMMs [196], or decision trees [197], and deep learning-based approaches [141, 198, 199, 200, 201, 59, 202, 203, 204]. [205, 163] showed that i-vectors are effective in predicting the severity of pathological speech. Motivated by the human ear, [163] introduced perceptually enhanced single frequency cepstral coefficients for pathological speech intelligibility measure. Pathological speakers often exhibit irregular rhythm patterns in their speech. To leverage this characteristic, [206, 207] explored rhythm-based features for severity assessment, utilizing SVMs and MLP classifiers. Their study highlighted the potential of rhythm as a key indicator of pathological speech impairment severity, though challenges remain in capturing the full complexity of rhythm variations across different speakers and integrating these features into more advanced machine learning models. Furthermore, patients with various pathological conditions often display distinct facial expressions. To harness these visual features, [208] introduced the first audio-visual pathological severity classification system using convolutional neural networks. [200] introduced a novel training framework, demonstrating that incorporating weak speech supervision can offer significant advantages in pathological speech severity classification. To leverage metadata information such as age, gender, and type of pathological condition, [209] proposed a multi-head attention-based multi-task framework. This approach jointly optimizes severity, type, gender, and age classifications for pathological speech, thereby enhancing the robustness of latent features across these additional factors. Recently, there has been a shift towards leveraging self-supervised speech embeddings, such as Whisper and wav2vec2, in pathological speech severity measurement. This approach is gaining popularity due to the limited availability of pathological data and the ability to utilize unlabeled and metadata from other datasets, making it a promising direction for improving severity assessment in resource-constrained scenarios [77, 47].
VIII Data Augmentation for
Pathological Speech Applications
To mitigate overfitting in deep learning models dealing with the low resource pathological speech data, data augmentation techniques have been applied for various tasks. Existing approaches to data augmentation in pathological speech can be broadly classified into the following categories: traditional augmentation, voice conversion, text-to-speech (TTS) synthesis, and generative models. Drawing inspiration from the temporal and speed differences between pathological and neurotypical speech, [210, 211, 84] proposed a data augmentation technique that involves adjusting the temporal and speed characteristics of neurotypical speech to simulate pathological speech. Similarly, [212] introduced a data augmentation technique that adjusts the phonetic-level tempo of healthy speech to resemble atypical speech, and vice versa. Their findings demonstrated that the former approach is more effective for pathological ASR. [213] employed a strategy that involved combining diverse pathological speech data from multiple languages to increase the number of data samples. Their results indicated that this approach offers certain advantages in pathological ASR. Additionally, they demonstrated that integrating data from various sources can improve model performance by providing a richer and more varied dataset for training. Generative models like GANs and variational autoencoders are also commonly employed to generate pathological speech samples from healthy speech [28, 173]. These synthetic samples serve dual purposes: they can be used for speech enhancement, improving the clarity and quality of pathological speech, and for expanding the dataset, thereby enhancing the diversity of training data. By increasing the variety of pathological speech examples, these models help in mitigating overfitting and improving the performance of deep learning models in tasks such as pathological speech detection and pathological ASR [28, 173, 214, 215]. Dysarthric speech datasets typically have a limited vocabulary, but using a conversion model can expand the set of out-of-vocabulary words, which can then be used for data augmentation. To this end, [78] utilized a voice conversion model to transform natural speech into synthetic pathological speech utterances.  [83] proposed a method incorporating temporal and spectral masking combined with time warping within a contrastive learning framework to develop robust speech representations for pathological ASR. [216] employed a TTS framework to generate synthetic dysarthric speech samples as an augmentation for pathological ASR. Although this approach showed some advantages, it faced challenges in generalizing to unseen speakers. This difficulty arose from the limited variability in pathological speech data, which was insufficient to fully capture the diverse characteristics of pathological speech.  [107] investigated different data augmentation techniques in pathological ASR and demonstrated that GAN-based perturbation methods are more effective than other augmentation approaches. However, a comprehensive investigation of data augmentation methods in pathological speech detection has not yet been conducted. It would be particularly important to explore various data augmentation techniques on clean datasets such as PC-GITA, rather than on the noisy datasets such as TORGO, which has been the case for the previously reviewed literature.
IX Other Challenges and Research Directions in Pathological Speech Processing
This section describes the various challenges and possible research directions for pathological speech analysis.
IX-A Robustness in Pathological Speech Detection
The methods developed so far have typically been designed and tested under specific environmental conditions and for a particular language. Each dataset used for pathological speech detection has its own unique variabilities in terms of both inter-speaker and intra-speaker differences. Additionally, different age groups exhibit distinct speaker attributes, adding complexity to pathological speech analysis. Language is another important factor; separate models have been independently developed for different languages. However, there is a need for a more robust model that is language-agnostic (to a possible degree given that pathological cues might be different in different languages), age-agnostic, accent-agnostic, and gender-agnostic to enhance overall effectiveness. Additionally, datasets such as Torgo are contaminated with noise. Researchers have shown that models trained on these datasets tend to learn environmental factors rather than focusing on genuine pathological cues [43]. To the best of our knowledge, the PC-GITA database is currently the only clean pathological dataset available under a research license. The development of pathological speech detection models robust to environmental distortions has been very limited, with only a few small studies addressing this area [217, 114, 113]. For example, [114] employed a test-time adaptation method to fine-tune pre-trained models on a validation set augmented with the test noise extracted from the test utterance. This method improves the robustness of state-of-the-art pathological speech detection methods, offering a promising solution to deploying such applications in realistic clinical settings. Similarly, [115] proposed an approach to resolve the noise disparity between the two groups of speakers in the TORGO database, such that models developed on this database learn pathology-discriminant cues instead of noise-discriminant ones. Besides robustness to environmental distortions, adversarial robustness of pathological speech detection models is another important topic and research direction. [115] explored the impact of acoustically imperceptible adversarial perturbations on deep learning-based pathological speech detection models. Results revealed a high vulnerability of such models to adversarial perturbations, with adversarial training ineffective in enhancing robustness.
IX-B Graph Neural Networks for Pathological Speech Detection
While many automatic pathological speech detection methods have demonstrated remarkable performance, the exploitation of common attributes such as pathological cues, speaker characteristics, age, and gender across different speakers remains limited. A promising direction would be to approach the pathological speech detection problem as a semi-supervised node graph classification task via graph neural networks (GNNs). This could involve constructing an inter-speaker graph from the utterances of various speakers, where the graph connectivity would help form speaker clusters based on the presence or absence of pathological cues. This structural information can then be leveraged through various graph message-passing mechanisms to improve detection accuracy and provide deeper insights into pathological speech patterns. Initial work in this direction has been done in [218], where a graph convolutional neural networks has been proposed for Parkinson’s disease detection method by framing the task as a node classification problem. In this approach, a graph is constructed using w2v2 speech embeddings across different speech segments from various patients as nodes, with the edges capturing the relationships between these segments. We believe that in this context, a promising research direction is to explore the incorporation of domain knowledge, such as establishing edges based on factors like gender, age, and severity scale of the patients.
IX-C Privacy
In pathological speech-based applications such as detection models or ASR systems, privacy-preserving solutions are crucial for safeguarding sensitive patient data. Since these systems process speech that reveals sensitive medical information, ensuring data security and confidentiality is of paramount importance. By employing privacy-preserving techniques - such as federated learning, differential privacy, and encryption - providers can ensure that individuals’ speech data is anonymized and never exposed to unauthorized parties. This not only fosters trust among patients but also complies with stringent data protection regulations, thereby mitigating the risk of breaches. Developing privacy-preserving pathological speech processing systems where one must balance two conflicting goals, i.e., increasing the utility of the models while preserving the privacy of the users, remains an important challenging research directions.
IX-D Multimodal Pathological Speech Analysis
Most existing systems primarily focus on leveraging speech cues for detecting pathological speech. However, complementary cues may also exist in visual forms, such as facial expressions or lip movements, which can provide additional insights. To effectively utilize these visual cues, multimodal frameworks like AV-HuBERT could prove advantageous, as they integrate both audio and visual information, potentially enhancing the accuracy and robustness of pathological speech detection models. Furthermore, combining other modalities, such as electroencephalogram (EEG) brain signals and textual data (transcripts), could offer a more complete characterization of disorders. These multimodal systems may help in capturing underlying neural, visual or linguistic patterns associated with speech pathology, improving detection and diagnosis across diverse patient groups.
X Explainability and Interpretibility
In the domain of pathological speech detection, explainability and interpretability are crucial due to their clinical significance. Although these terms are often used interchangeably because of their close overlap, it’s important to distinguish between them. As defined by [219], interpretable models are designed to be inherently understandable, whereas explainable models provide post-hoc explanations for the decisions made by existing ”black box” systems that are otherwise incomprehensible to humans. [220] further clarified the distinction by defining interpretability as ”How does the model work” and explainability as ”What else can the model tell me.” Keeping these distinctions in mind, particularly in the context of pathological speech detection, there has been very limited, if any, attention given to this aspect of research [221, 222, 223, 224]. Most existing studies focus on improving model accuracy and performance, often overlooking the importance of making these models interpretable and explainable. This gap is critical, as the ability to understand and explain the decision-making process in pathological speech detection models can provide valuable insights for clinicians and support more trustworthy and reliable use of artificial intelligence in healthcare applications. To address this issue, [221] attempted to map high-dimensional acoustic features to sound pattern of English binary phonological features, recognizing that these high-dimensional acoustic features are challenging to interpret in clinical practice. The problem with this approach is that it may be biased toward English sounds, making it susceptible to errors when applied to different phonetic environments. Further, [224] applied the SHAP algorithm to identify the most influential features in their model and discovered that constant-vowel transitions were the key features in reversing the model’s decisions in pathological speech detection. Recently, [225] demonstrated using canonical correlation analysis that the human frequency range (0–210 Hz) plays a critical role in influencing model decisions for pathological speech detection. Although some research has been conducted, substantially more research is needed to address this under-explored but very important area.
XI Large Language Models
for Pathological Speech
Given the recent advancements in large language models (LLMs), which have demonstrated significant progress across a wide range of applications [226, 227], exploring their potential for pathological speech analysis appears to be an inevitable and promising direction. LLMs could be leveraged for tasks such as detecting speech impairments based on specific phrase cues, improving the automatic classification of speech disorders, and enhancing the reconstruction of unintelligible or difficult speech. Additionally, LLMs’ ability to capture complex linguistic patterns may lead to more accurate models for personalized therapy, rehabilitation, and assistive communication tools for individuals with speech impairments. This exploration could bridge the gap between traditional speech processing techniques and state-of-the-art language models, opening up new avenues for more effective and adaptive speech rehabilitation systems.
XII Conclusion
This paper provides a comprehensive overview of deep learning applications for pathological speech, encompassing detection, recognition, intelligibility assessment and enhancement. Additionally, it compiles a thorough list of both accessible and non accessible pathological speech datasets, which will serve as valuable resources for future research and accelerate progress in the field. To the best of our knowledge, this is the first review to offer an in-depth survey of pathological speech from the perspectives of detection, recognition, enhancement, explainibility and data augmentation. Furthermore, it outlines potential future research directions, particularly in the context of robust and interpretable models that can be deployed in clinical practice.
Acknowledgment
This work was supported by the Swiss National Science Foundation project CRSII5_202228 on “Characterisation of motor speech disorders and processes”.

### v2

@@
-Deep Learning for Pathological Speech: A Survey
+Overview of Automatic Speech Analysis and Technologies for Neurodegenerative Disorders: Diagnosis and Assistive Applications
 Abstract
@@
 It also highlights key challenges, such as ensuring robustness, privacy, and interpretability.
-The paper concludes by exploring promising future directions, including the adoption of multimodal approaches and the integration of graph neural networks and large language models to further advance speech technology for neurodegenerative speech disorders.
+The paper concludes by exploring promising future directions, including the adoption of multimodal approaches and the integration of large language models to further advance speech technologies for neurodegenerative speech disorders.
 Index Terms:
@@
 I Introduction
-Speech production is a complex mechanism involving cognitive planning, coordinated muscle activity, and sound creation [1].
-The process starts in the brain with the conceptualization of a message, followed by the organization of phonetic and prosodic plans such as rhythm and style.
-The motor cortex then orchestrates the activation of approximately $100100$ muscles, enabling the articulatory organs such as the tongue, lips, and jaw, to shape the vocal tract and produce specific sounds.
-The initial sound is generated in the larynx, where air from the lungs causes the vocal folds to vibrate.
+Speech production is a complex mechanism that involves cognitive planning, coordinated muscle activity, and sound creation [1].
+The process starts in the brain with the conceptualization of a message, followed by the organization of phonetic and prosodic plans, such as rhythm and style.
+The motor cortex then orchestrates the activation of approximately $100$ muscles, allowing articulatory organs such as the tongue, lips, and jaw to shape the vocal tract and produce specific sounds.
+The initial sound is generated in the larynx, where the air from the lungs causes the vocal folds to vibrate.
 These phonatory structures adjust voice quality and prosody, while the articulatory organs further refine the sound by altering the shape of the vocal tract.
 Finally, the resulting speech is emitted through the oral and nasal cavities.
-Given the intricate coordination required for this process, any disruption to these finely tuned mechanisms can severely impair communication.
-Neurodegenerative disorders such as Parksinson’s disease (PD), Amyotrophic Lateral Sclerosis (ALS), or Alzheimer’s disease are leading causes of such disruptions [2, 3, 4].
-These disorders impair the brain regions and motor systems responsible for initiating, planning, and controlling the movements needed for speech production, resulting in a variety of speech disorders such as dysarthria, aphasia, apraxia of speech, dysphonia, or stuttering [5, 6, 7, 4, 8, 9].
-Dysarthria and apraxia of speech, commonly seen in PD and ALS, are characterized by articulation deficiencies, vowel distortions, reduced loudness variation, hypernasality, or syllabification [5, 4].
-Dysphonia, also frequent in PD and ALS, is marked by abnormal voice quality such as hoarseness and breathiness [6].
-In contrast, aphasia typically presents as difficulties with word-finding, sentence construction, and understanding spoken or written language, and is most commonly associated with Alzheimer’s disease or other forms of dementia [8].
-Lastly, stuttering, characterized by disruptions in the flow of speech such as repetitions or prolongations can also arise in neurodegenerative disorders like Parkinson’s disease [9].
-Because of increasing population numbers and aging, the prevalence of neurological disorders, and hence, of these various speech disorders, is also rapidly increasing.
-In $20192019$, the World Health Organization estimated that over $8.58.5$ million people worldwide were living with PD [10], up from $6.16.1$ million in $20162016$ [11] and $2.52.5$ million in $19901990$ [11].
-By $20402040$, this figure is projected to surpass $1717$ million [11].
-Similarly, dementia affected more than $4646$ million people globally in $20152015$ and this figure is expected to rise to $131.5131.5$ million by $20502050$ [12].
-The prevalence of ALS is also growing significantly, with cases anticipated to increase by nearly 70% between $20152015$ and $20402040$ [13].
+Given the intricate coordination required for this process, any disruption of these finely tuned mechanisms can severely alter communication and result in pathological speech.
+In this work, we define pathological speech as speech that deviates from neurotypical patterns due to underlying impairments.
+These deviations can manifest along multiple dimensions, including voice, articulation, prosody, and language.
+Voice impairments involve abnormalities in vocal fold vibration or in breath control, leading to hoarseness, breathiness, or a strained voice [2, 3].
+Articulation impairments involve abnormalities in the coordination or movement of the various articulators, leading to slurred, imprecise, or segmented speech [4, 5].
+Furthermore, prosodic impairments involve abnormalities in the rhythm, stress, or intonation of speech, leading to speech that may sound flat or monotonic [6].
+Finally, language impairments involve abnormalities in the formulation or comprehension of linguistic content, leading to difficulties with word retrieval, sentence construction, or understanding spoken or written language [7].
+While pathological speech can arise from a wide range of neurological, structural, or functional impairments, our objective is to provide an overview of automated methods and speech-based technologies targeting pathological speech arising due to neurodegenerative disorders.
+Neurodegenerative disorders such as Parkinson’s disease (PD), Amyotrophic Lateral Sclerosis (ALS), or Alzheimer’s disease are leading causes of voice, articulation, prosody, and language disruptions [8, 9, 4].
+These disorders impair the brain regions and motor systems responsible for initiating, planning, and controlling the movements needed for speech production, resulting in a variety of speech disorders such as dysarthria, aphasia, apraxia of speech, or dysphonia [5, 2, 10, 4, 7, 11, 12].
+Dysarthria and apraxia of speech, commonly seen in PD and ALS, are primarily characterized by articulatory and prosodic deficiencies [13], vowel distortions, reduced loudness variation, hypernasality, or syllabification [5, 4].
+Dysphonia, also frequent in PD and ALS, is marked by abnormal voice quality such as hoarseness and breathiness [2, 3].
+In contrast, aphasia typically presents as difficulties with word-finding and sentence construction, and is most commonly associated with Alzheimer’s disease or other forms of dementia [7].
+As the population grows and ages, the prevalence of neurological disorders, and consequently of various speech disorders, rapidly increases.
+In $2019$, the World Health Organization estimated that over $8.5$ million people worldwide were living with PD [14], up from $6.1$ million in $2016$ [15] and $2.5$ million in $1990$ [15].
+By $2040$, this figure is projected to surpass $17$ million [15].
+Similarly, dementia affected more than $46$ million people globally in $2015$ and this figure is expected to rise to $131.5$ million by $2050$ [16].
+The prevalence of ALS is also growing significantly, with cases anticipated to increase by nearly 70% between $2015$ and $2040$ [17].
 This increasing prevalence of neurological disorders, and consequently of the associated speech disorders, underscores the need to prioritize speech disorders both in the context of clinical practice as well as in the context of speech-based technologies.
-In clinical practice, diagnosing the presence of a speech disorder (i.e., discriminating between neurotypical and impaired speech) is crucial, since the presence of a disorder can be one of the earliest signs of neurodegenerative disorders [14, 15, 16].
-Further, an accurate differential diagnosis of the speech disorder (e.g., discriminating between dysarthria and apraxia of speech) is also important, since it can provide clues about the underlying neuropathology [17, 18].
-Additionally, monitoring speech characteristics such as severitys and intelligibility after diagnosis is essential for tracking disease progression and assessing the effectiveness of speech therapy interventions over time [2, 19].
-Traditionally, clinicians conduct such evaluations through costly and time-consuming auditory-perceptual assessments as depicted in Fig.
-2.
-However, because of the difficulty of detecting clinical-perceptual characteristics by ear (particularly in the presence of mild impairments) and because the clinical-perceptual characteristics of several speech disorders (e.g., of dysarthria and apraxia of speech) overlap, diagnosing the presence of a disorder and discriminating between various disorders is hard for non-expert clinicians and even expert inter-rater agreement can be low [20, 21].
-In the context of speech-based technologies, patients with speech disorders often struggle to access and effectively use various virtual assistant tools like Cortana, Alexa, and Siri [22].
+Accurately diagnosing the presence of speech disorders in clinical practice (i.e., distinguishing between neurotypical and impaired speech) is crucial, as the presence of such disorders may serve as an early indicator of neurodegenerative conditions [18, 19, 20].
+Further, an accurate differential diagnosis (e.g., discriminating between dysarthria and apraxia of speech) can provide important clues about the underlying neuropathology [21, 22].
+Monitoring speech characteristics such as severity and intelligibility after diagnosis is also essential for tracking disease progression and evaluating the effectiveness of speech therapy interventions over time [8, 23, 12, 6].
+Speech assessment in clinical practice relies on established perceptual evaluation scales that serve as gold standards for diagnosing and characterizing various aspects of speech impairments.
+For example, the GRBAS scale [24, 25] evaluates voice quality through parameters like grade, roughness, breathiness, asthenia, and strain.
+The Consensus Auditory Perceptual Evaluation-Voice (CAPE-V) [23] offers a similar perceptual framework, excluding asthenia.
+In addition to these general assessment scales, specific scales are employed for specific conditions.
+For instance, the Unified Parkinson’s Disease Rating Scale (UPDRS) [26] includes components for speech and motor function assessment in PD, while the Bogenhausen Dysarthria Scales (BoDyS) [27] focus on the severity and profile of dysarthric impairments.
+Intelligibility, a key outcome measure in many disorders, is often assessed using tools such as the Assessment of Intelligibility of Dysarthric Speech [28], which standardizes both single-word and sentence intelligibility evaluations.
+Traditionally, clinicians conduct these evaluations through costly and time-consuming auditory-perceptual assessments [12, 23], as illustrated in Fig.1.
+Diagnosing speech disorders and distinguishing between various conditions can be particularly challenging, even for experienced clinicians.
+This difficulty arises from the subtle nature of clinical-perceptual characteristics which are often hard to detect by ear, especially in cases of mild impairments.
+The overlapping characteristics of certain speech disorders, such as dysarthria and apraxia of speech, further complicate the process.
+Consequently, inter-rater agreement for (differential) diagnosis of speech disorders among clinicians can be low [29, 30].
+These clinical challenges highlight the growing need for complementary, technology-driven approaches to support diagnosis, monitoring, and intervention.
+In parallel, the increasing prevalence of speech disorders poses significant accessibility challenges in patients’ everyday interactions with speech-based technologies.
+For example, individuals with dysarthria or apraxia of speech often experience difficulty using mainstream virtual assistants such as Cortana, Alexa, and Siri [31].
 As speech disorders become more prevalent with neurodegenerative conditions, it is critical to prioritize both the diagnosis and treatment of these disorders in clinical practice, while also ensuring that patients with such impairments have equitable access to speech-based technologies.
 Addressing these barriers could lessen the burden on the health care system and significantly improve the patients’ quality of life and their ability to engage with everyday digital tools.
-Figure 1: Traditional auditory-perceptual assessment in clinical practice. The clinician listens to the (potential) patient speaker and assesses by ear the various characteristics of the speech.
-Figure 2: Automatic pathological speech analysis system.
+Figure 1: Traditional auditory-perceptual assessment in clinical practice (bounded by the dashed box A) and automatic pathological speech analysis system (bounded by the solid box B).
+The clinician listens to the (potential) patient and assesses by ear the various characteristics of the speech.
 The automatic model is trained to detect and analyze speech impairments.
@@
 Additionally, they may perform manual acoustic analysis to gain further insights into the patient’s speech patterns and impairments, providing complementary information to the automatic model’s decision.
-Aiming to assist the clinical diagnosis and treatment of patients suffering from neurodegenerative disorders, there has been a growing interest in the research community to develop automated and objective methods for pathological speech analysis as shown in Figure 2.
+Aiming to assist the clinical diagnosis and treatment of patients suffering from neurodegenerative disorders, there has been a growing interest in the research community to develop automated and objective methods for pathological speech analysis.
+A schematic illustration of such analysis is depicted in Fig.
+1.
 These advanced technologies are designed to minimize bias, enhance diagnostic accuracy, and streamline the assessment process, ensuring greater efficiency and consistency.
@@
 Besides the clinical domain, efforts have been directed towards developing various speech-based technological applications aimed at pathological speakers, such as automatic speech recognition systems (ASRs) [32], speech synthesis systems [33, 34, 35, 36] or intelligibility enhancement solutions [37, 38, 39, 40].
-To the best of our knowledge, there is currently no comprehensive survey paper discussing the research directions and challenges of this area both from a clinical as well as a technological perspective.
-Although there have been related works such as [32], they primarily provide an overview of acoustic-articulatory features in neurodegenerative disorders.
-Other studies such as [33, 34, 35, 36, 37] mainly focus on the discrimination between neurotypical and impaired speech.
-In $20162016$, [38] discussed some of the wider challenges faced in the pathological speech domain.
-However, this work addresses only a limited set of challenges and is now somewhat outdated, given the rapid advancements in the field over the past decade.
-In contrast, as depicted in Figure 3, our work aims to fill this gap by providing an extensive review of the field encompassing pathological speech from various clinical and technological perspectives, such as automatic discrimination between neurotypical speech and speech disorders, ASR systems for pathological target speakers, enhancement systems aiming to enhance the intelligibility of pathological speech, severity and intelligibility assessment and data augmentation approaches.
-Figure 3: Overview of Deep Learning for Pathological Speech: This framework leverages various input features and datasets processed through deep learning models for tasks including detection, recognition, enhancement, assessment and explainaibility of pathological speech.
-It employs a range of metrics to evaluate model performance, providing a comprehensive approach to improving speech impairment detection and treatment.
+To our knowledge, there is currently no comprehensive survey paper that discusses the research directions and challenges of this area both from a clinical and a technological perspective.
+Although, there have been related works, such as [41, 42, 43], they are primarily focused on specific characteristics, applications, or disorders, limiting their scope.
+For instance, [44] provides an overview of acoustic-articulatory characteristics in neurodegenerative disorders.
+Other studies, such as [45, 46, 47, 48, 49, 50], mainly focus their review on the discrimination between neurotypical and impaired speech.
+Gupta et al.
+[51] discussed some of the wider challenges faced in the pathological speech domain.
+However, this work addresses only a limited set of challenges and is now somewhat outdated given the rapid advancements in the field over the past decade.
+In contrast, as depicted in Fig.
+2, our work aims to fill this gap by providing an extensive review of the field encompassing pathological speech from various clinical and technological perspectives, such as automatic discrimination between neurotypical speech and speech disorders, ASR systems for pathological target speakers, enhancement systems aiming to enhance the intelligibility of pathological speech, severity and intelligibility assessment, and data augmentation approaches.
+Figure 2: Overview of key components discussed in this manuscript, including datasets, features, and research directions for pathological speech.
 The remainder of the paper is organized as follows.
-Section II provides a high level overview of the handcrafted acoustic features used in pathological speech analysis.
-Section III describes various accessible and non accessible pathological datasets employed in the literature.
+Section II describes various pathological datasets employed in the literature.
+Section III provides a high level overview of the various speech representations used in pathological speech analysis.
 Section IV describes the different approaches to automatic pathological speech detection.
@@
 Section VI discusses speech enhancement techniques aimed at improving pathological speech intelligibility.
-Section VII describes approaches for automatically estimating the severity and intelligibility of pathological speakers.
+Section VII describes approaches for automatically estimating the intelligibility and severity of pathological speakers.
 Section VIII summarizes the various data augmentation methods used in automatic pathological speech systems.
-The robustness issues arising in this field are discussed in Section IX-A.
-Finally, Section X explores the interpretability of pathological speech models.
+Finally, Section IX presents challenges and promising future research directions in the field.
 In summary, our contributions are the following:
   1. 1.
-    We present the first comprehensive survey of automatic approaches for pathological speech from a clinical and technological perspectives ranging from detection, recognition, enhancement, assessment and interpretability.
+    We present the first comprehensive survey of automatic approaches for pathological speech from a clinical and technological perspectives ranging from detection, recognition, enhancement, and assessment.
   2. 2.
     We highlight current limitations in the field and propose several promising future directions for research in pathological speech processing.
-II Pathological Speech: Acoustic Characteristics and Representation
-Pathological speech exhibits a range of acoustic anomalies, including deviations in pitch, loudness, vowel space reduction, and articulation [39, 40].
-Additionally, it can lead to asymmetrical tension in the vocal folds, resulting in irregular vibrations and, consequently, an abnormal fundamental frequency.
-Excessively high or low fundamental frequency, combined with excessive vocal intensity, can exacerbate the severity of pathological voice conditions, producing characteristics such as shrillness, screechiness, hoarseness, or huskiness [41].
-Moreover, patients (such as those diagnosed with PD) can also show inconsistent rhythmic structures in comparison to control groups under various auditory feedback conditions [42]
-To capture such abnormalities as biomarkers for automatic pathological speech detection, researchers have employed various handcrafted acoustic features such as OpenSMILE [43], Mel-frequency cepstral coefficients (MFCCs) [44], or spectro-temporal sparsity features [45].
-Further, various time-frequency representations such as the short-time Fourier transform (STFT) or the Mel spectrogram have also been directly exploited in combination with deep learning approaches to directly learn pathology-discriminant cues [46].
-More recently, latent embeddings derived from self-supervised models have been used as more powerful representations of speech patterns and the various impairments [43, 47, 48].
-It is important to note that a significantly larger set of features and representations have been employed to characterize disorders across various pathologies than the ones outlined above.
-However, here we focus on highlighting the most commonly utilized ones, which will be further elaborated upon in the following.
-II-A Handcrafted Acoustic Features
-OpenSMILE
-Among the various features used in pathological speech detection, OpenSMILE features have been widely explored in the literature in conjunction with traditional machine learning algorithms [49, 50, 51, 52, 31, 53, 54, 55, 56, 57, 58, 59, 60, 61].
-The OpenSMILE feature set includes a 6552-dimensional feature vector that primarily consists of low-level audio features such as CHROMA, CENS, loudness, MFCCs, and other spectral features.
-In the context of pathological speech, these features can capture the subtle anomalies in speech patterns that are indicative of disorders.
-OpenSMILE’s comprehensive feature set allows for the analysis of prosodic elements like pitch, jitter, shimmer, and formant frequencies, which are crucial for identifying and differentiating various speech pathologies.
-However, OpenSMILE features are very general features that have been used for a variety of speech applications such as emotion recognition [62], speech recognition [63] and they are not specifically handcrafted to capture pathological cues, thus may not be the most suitable features for pathological speech analysis.
-MFCCs
-MFCCs have garnered significant attention in pathological speech detection due to their potential to characterize articulation deficiencies [64, 65, 66, 67, 68, 69, 70, 71, 72].
-While MFCCs have demonstrated promising results, they exhibit high sensitivity to background noise and present challenges in real-time applications due to their complex nature and lack of interpretability.
-Spectro-temporal sparsity
-Since pathological speech can be breathy, semi-whispery, and is characterized by abnormal pauses and imprecise articulation, it can be expected that its spectro-temporal sparsity differs from the spectro-temporal sparsity of neurotypical speech.
-To characterize spectro-temporal sparsity, various sparsity-based features have been introduced in [45, 73].
-Although such features have been shown to be discriminative of various speech disorders[45, 73, 61, 74], they are highly sensitive to environmental artefacts such as noise and reverberation.
-II-B Time-Frequency Representations
-Input representations such as the STFT and its variants allow for the analysis of speech signals in both time and frequency domains, providing valuable insights into speech characteristics such as pitch, formant shifts, and spectral irregularities.
-The STFT is often employed due to its ability to capture dynamic changes in the signal over time, which is critical for identifying variations in speech patterns linked to speech disorders [75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 25, 85, 86].
-Furthermore, wavelet transforms and the Continuous Wavelet Transform offer better time-frequency localization, which is especially beneficial for analyzing transient and non-stationary features of pathological speech [87].
-II-C Self Supervised Embeddings
-Even though handcrafted acoustic features and time-frequency input representations have achieved promising results in the analysis of pathological speech, their performance remains limited.
-The current state-of-the-art input representations for pathological speech primarily consist of latent embeddings derived from self-supervised models such as wav2vec 2.0 (w2v2) [88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 47, 101, 102, 103, 104, 61, 105, 106, 107, 108, 109, 110, 111].
-These advanced models have proven to be more effective in capturing complex speech patterns and nuances, leading to improved performance in tasks like detection and recognition [112, 48], However, there is still ongoing research to further refine these embeddings and explore additional models to enhance their robustness and generalization across diverse pathological speech conditions [113, 114, 115].
-III Pathological Speech Datasets
-Datasets for pathological speech research remain scarce due to several inherent challenges in data collection.
+II Pathological Speech Datasets
+Datasets for pathological speech research are scarce due to several inherent challenges in data collection.
 One of the primary difficulties is the sensitive nature of the population involved.
@@
 Moreover, speech impairments may fluctuate over time, further complicating the process of capturing representative speech samples.
-When collecting datasets for pathological speech research, it is essential to prioritize inclusivity, ensuring a diverse representation of different speech disorders, age groups, and demographics.
+When collecting datasets for pathological speech research, it is essential to prioritize inclusivity, i.e., ensuring a diverse representation of different speech disorders, age groups, and demographics.
 Additionally, data collection protocols should consider the comfort and cooperation of participants.
@@
 Lastly, it is important to design flexible and scalable collection methods that can capture a range of speech characteristics while maintaining consistent quality across diverse individuals and conditions.
-In the remainder of this section, we briefly review pathological datasets commonly used in the literature.
-Please note that due to restricted access to most of the pathological neurodegenrative datasets, we are unable to provide detailed information such as duration, number of male and female speakers, etc., for all of them.
-A summary of these datasets and their characteristics is presented in Table I.
+In the remainder of this section, we briefly review pathological datasets [^1] Due to restricted access to many datasets, we are unable to provide complete metadata, including details like the number of male and female speakers commonly used in the literature.
+The summary of these datasets and their characteristics is presented in Table I.
   • TORGO  [52].
-The TORGO dataset contains speech recordings from English speaking subjects (controls and patients) and the corresponding three dimensional articulatory features.
+The TORGO dataset contains English (spontaneous and read) speech recordings from control speakers and patients and the corresponding three-dimensional electromagnetic articulography (EMA).
 The patients suffer from ALS or Cerebral Palsy (CP).
-The dataset consists of seven (three females, four males) controls and eight (three females, five males) dysarthric speakers respectively.
-PC-GITA [117].
-The PC-GITA dataset is a Spanish dataset containing speech recordings from $5050$ controls and $5050$ patients suffering from PD.
-The two groups of speakers are age- and gender-matched, with $2525$ female and $2525$ male speakers in each group.
-MoSpeeDi [118].
-The MoSpeeDi dataset is a French dataset containing speech recordings from $466466$ controls and $138138$ patients suffering from various types of motor speech disorders such as dysarthria or apraxia of speech.
-While subgroups of age- and gender-matched controls and patients can be found within the database, the overall dataset is not age- and gender-matched.
-Nemours [119].
-The Nemours dataset is an English dataset containing speech recordings from $1111$ male speakers with varying degrees of dysarthria severity.
-This dataset includes a total of 74 sentences and 814 short nonsense sentences.
-CUDYS [120].
-The Cantonese Dysarthric Speech Corpus (CUDYS) contains speech recordings from $55$ controls and $1111$ dysarthric patients.
-The control group consists of $22$ female and $33$ male speakers whereas the dysarthric group consists of $55$ female and $66$ male speakers.
-The dataset has a total duration of $1010$ hours, with the speakers’ ages ranging between $2525$ and $6767$ years.
-AMSDC [121].
-The Atlanta Motor Speech Disorders Corpus (AMSDC) is an English dataset containing speech recordings from $9999$ ($6262$ males and $3737$ females) patients diagnosed with neurogenic disorders.
-This dataset is essential for studying motor speech disorders, though access to the full dataset might be restricted.
-EST [122].
-The EST dataset is a Dutch dataset containing speech recordings from 16 dysarthric patients ($1616$ males).
-The dataset has a total duration of $\approx$ $66$ hours and the speakers’ ages range between $3434$ and $7575$ years.
-EasyCall [123].
-The EasyCall dataset is an Italian dataset containing speech recordings from $2424$ controls and $3131$ patients diagnosed with PD, Huntington’s disease, ALS, and peripheral neuropathy.
-The control group consists of $1010$ female and $1414$ male speakers whereas the patient group consists of $1111$ female and $2020$ male speakers.
-The corpus includes a total of 21,386 speech commands, with 10,077 recordings from healthy subjects and 11,309 from dysarthric patients.
-COPAS [124].
-The Corpus of Pathological and Normal Speech (COPAS) contains speech recordings from Dutch speakers with $197197$ pathological speakers and $122122$ normal speakers.
-The database comprised of 8 distinct pathological categories.
-ParkCeleb [125].
-The previously reviewed datasets are not longitudinal and do not allow tracking the progression of the speech disorder within the same patient along time.
-To address this gap, the English ParkCeleb dataset was recently introduced in [125].
-This dataset contains speech recordings from $4040$ controls and $4040$ patients suffering from PD.
-Summary of datasets.
-As outlined in Table I, many of the pathological speech datasets used in the literature are private and not openly accessible, limiting their availability for broader research and replication.
-Furthermore, datasets such as TORGO, although publicly available, have been shown to contain detrimental recording artefacts [43], which limits their suitability for studying pathological speech characteristics or for developing automatic pathological speech detection approaches [43, 113].
-This lack of accessible, high-quality data creates a substantial barrier for developing and validating effective algorithms and tools.
-Consequently, the advancement in understanding and treating pathological speech conditions is slowed, limiting the potential for innovative solutions and breakthroughs in this critical area of healthcare.
-Ensuring the availability of clean, reliable, and comprehensive datasets is essential to foster meaningful research and development.
-Additionally, many existing pathological speech datasets are predominantly in English, restricting their applicability to non-English-speaking populations and limiting cross-linguistic comparisons.
-This linguistic bias creates challenges when attempting to develop universal speech models or tools that can be applied globally.
-Furthermore, most of these datasets are relatively small, often comprising only a limited number of participants, which hinders the development of robust, generalizable models.
-The small size of these datasets also makes it difficult to account for the variability in speech disorders across different populations and conditions.
-Another important limitation is the imbalance in demographic representation in terms of gender.
-Many datasets suffer from a skewed representation, often over-representing one gender, which can lead to biases in speech models.
-Additionally, some datasets may not include a sufficient variety of age groups or pathological conditions, further restricting their generalizability.
-These factors collectively underscore the need for high quality, larger, more diverse, and well-balanced pathological speech datasets that reflect the complexity and diversity of real-world populations.
-Figure 4: Automatic Pathological Speech Detection Pipeline: The process begins with recording the patient’s speech, which is then processed to extract features that are meaningful and suitable for an AI classifier.
-These features are transformed into a feature space, where a neural network model analyzes them to make a diagnostic decision.
-This pipeline streamlines the detection process, enabling efficient and accurate identification of pathological speech patterns.
-IV Automatic Pathological Speech Detection
-TABLE I: Overview of Pathological Speech Databases.
-(PD: Parkinson’s disease, CVA: Cerebrovascular accident, TBI: Traumatic brain injuries, ALS: Amyotrophic lateral sclerosis, Huntington’s disease, 3D: Articulatory (3D), L: Longitudinal data, Accessible: Accessible through a license for research purposes )
+The dataset consists of recordings from 77 (33 female, $4$ male) control speakers and 88 (33 female, $5$ male) patients.
+    TABLE I: Overview of pathological speech datasets (PD: Parkinson’s disease, CVA: Cerebrovascular accident, TBI: Traumatic brain injuries, ALS: Amyotrophic lateral sclerosis, Huntington’s disease, AD: Alzheimer’s Disease, MCI: Mild Cognitive Impairment, EMA: Electromagnetic Articulography, EGG: Electroglottography, L: Longitudinal data).
+Public datasets are openly available on the web and can be downloaded without the need for explicit approval or signing an agreement.
+Accessible datasets require an application process or signing an agreement, but can be obtained by the research community under defined conditions.
+Non-accessible datasets are private and not available to the wider research community.
     Database Number of Speakers Type of Impairment Language Modality Accessibility
@@
     PC-GITA 50 50 PD Spanish Speech Accessible
-TORGO 07 08 ALS/CP English Speech + (3D) Public
+    TORGO 07 08 ALS/CP English Speech + EMA Public
     MoSpeeDi 466 138 Dysarthira/Apraxia French Speech Not Accessible
@@
     EasyCall 24 31 PD, HD, ALS Italian Speech Public
+    Saarbrucken Voice Database 869 1,356 Pathological German Speech + EGG Public
+    Turkish Parkinson Dataset 20 20 PD Turkish Speech Not Accessible
+    Czech Parkinson’s Dataset 22 61 PD Czech Speech Public
+    Italian Parkinson’s Database 28 37 PD Italian Speech Public
+    NeuroVoz 58 54 PD Spanish Speech Public
     COPAS 197 122 Dysarthira and others Dutch Speech Accessible
-ParkCeleb 40 40 PD English Speech (L) Accessible
-Neurodegenerative diseases such as PD and ALS can significantly impact various sensorimotor and articulatory processes, leading to pathological speech impairments such as dysarthria, apraxia of speech, aphasia, dysphonia.
-As discussed in Section I, clinicians traditionally rely on laborious and time-consuming auditory-perceptual measures to diagnose these speech impairments accurately.
+    EWA-DB 896 226 PD, AD, MCI Slovak Speech Accessible
+    ParkCeleb 40 40 PD English Speech + Video + L Accessible
+    mPower 5,581 1,087 PD English Multimodal Accessible
+  • PC-GITA [53].
+The PC-GITA dataset contains Spanish (spontaneous and read) speech recordings from $50$ control speakers and $50$ patients suffering from PD.
+The two groups of speakers are age- and sex-matched, with $25$ male and $25$ female speakers in each group.
+  • MoSpeeDi [54].
+The MoSpeeDi dataset contains French (spontaneous and read) speech recordings from $466$ control speakers and $138$ patients suffering from various types of motor speech disorders such as dysarthria or apraxia of speech.
+While subgroups of age- and sex-matched controls and patients can be found within the database, the overall dataset is not age- and sex-matched.
+  • Nemours [55].  The Nemours dataset is an English dataset containing (read) speech recordings from $11$ male speakers with varying degrees of dysarthria severity.
+  • CUDYS [56].
+The Cantonese Dysarthric Speech Corpus contains Cantonese (read) speech recordings from $5$ control speakers and $11$ patients suffering from cerebellar degeneration.
+The control group consists of $2$ female and $3$ male speakers, whereas the dysarthric group consists of $5$ female and $6$ male speakers.
+  • AMSDC [57].
+The Atlanta Motor Speech Disorders Corpus is an English dataset containing (spontaneous and read) speech recordings from $99$ ($62$ male, $37$ female) patients diagnosed with various disorders such as PD, ALS, or dementia.
+  • EST [58].  The EST dataset is a Dutch dataset containing (read) speech recordings from $16$ male dysarthric patients due to PD, traumatic brain injuries, or cerebrovascular accident. .
+  • EasyCall [59].
+The EasyCall dataset is an Italian dataset containing (read) speech recordings from $24$ controls and $31$ patients diagnosed with PD, Huntington’s disease, ALS, and peripheral neuropathy.
+The control group consists of $10$ female and $14$ male speakers, whereas the patient group consists of $11$ female and $20$ male speakers.
+  • COPAS [60].
+The Corpus of Pathological and Normal Speech dataset is a Dutch dataset containing (spontaneous and read) speech recordings from $197$ pathological speakers and $122$ control speakers.
+The database comprises $8$ distinct pathological categories such as dysarthria, hearing impairment, cleft, etc.
+  • ParkCeleb [61].
+The previously reviewed datasets are not longitudinal and do not allow tracking the progression of the speech disorder within the same patient along time.
+To address this gap, the English ParkCeleb dataset was recently introduced in [61].
+This dataset contains (spontaneous) audio-visual recordings (such as studio interviews or press conferences) from $40$ (22 female, $38$ male) control speakers and $40$ (22 female, $38$ male) patients suffering from PD.
+  • Italian Parkinson’s Database [62].
+The dataset contains Italian (read) speech recordings from a total of $65$ speakers, including $28$ patients with PD and $37$ control speakers.
+Sex distribution is non-uniform across groups, with the PD group consisting of $19$ male and $9$ female speakers and the control group consisting of $23$ male and $14$ female speakers.
+  • NeuroVoz [63].
+The NeuroVoz dataset contains ( spontaneous, read, and listen and repeat) speech recordings from $112$ Castilian Spanish speakers, including $54$ patients diagnosed with PD and $58$ control speakers.
+The control group consists of $28$ male speakers, $26$ female speakers, and $1$ speaker whose sex information is not provided, whereas the patient group consists of $33$ male and $20$ female speakers.
+  • Saarbrücken Voice Database [64] [^2] https://stimmdb.coli.uni-saarland.de/ (accessed July 07, 2025).
+The Saarbrücken Voice Database contains (read) speech recordings and electroglottography data from $2,255$ German speakers, including $1,356$ patients and $869$ control speakers [65].
+Patients are diagnosed with various voice disorders.
+The control group consists of $433$ male and $436$ female speakers, whereas the patient group consists of $629$ male and $727$ female speakers.
+  • Turkish Parkinson Speech Dataset [66].
+This dataset contains (read) speech recordings from $40$ Turkish speakers, including $20$ controls and $20$ patients with PD.
+The PD group contains $6$ female and $14$ male speakers, whereas the control group contains $10$ female and $10$ male speakers.
+  • Czech Parkinson’s Dataset [67].
+This dataset contains (read) speech recordings from $83$ Czech speakers, including $22$ control speakers and $61$ patients diagnosed with PD or atypical parkinsonian syndromes.
+The patient group consists of $30$ female and $31$ male speakers, whereas the control group consists of $11$ female and $11$ male speakers.
+  • EWA-DB [68].
+The EWA-DB dataset consists of (spontaneous and read) speech recordings of $1,122$ Slovak speakers, including $896$ control speakers and $226$ patients diagnosed with PD, mild cognitive impairment, or Alzheimer’s disease.
+The patient group consists of $121$ male and $105$ female speakers, whereas the control group consists of $248$ male and $648$ female speakers.
+  • mPower Parkinson’s Dataset [69].
+The mPower study collected multimodal smartphone sensor data from $6805$ participants, including $1,087$ patients diagnosed with PD and $5,581$ control speakers.
+Patient and control status were self-reported.
+The dataset was collected through the mPower mobile application and includes four distinct sensor-based assessment modalities, i.e., spatial memory evaluation, gait analysis through walking tasks, manual dexterity measurement via finger tapping, and vocal function assessment using sustained phonation recordings.
+As presented in Table I, a considerable number of the datasets used in the literature are private and not openly accessible.
+This limits their availability for broader research and replication efforts.
+Although some datasets such as TORGO are publicly available, they suffer from detrimental recording artifacts [70], which compromise their usefulness for studying pathological speech characteristics and for developing reliable automatic detection methods [70, 71].
+The lack of accessible, high-quality data poses a major challenge to the development and validation of effective algorithms, slowing progress in understanding and addressing pathological speech conditions.
+Ensuring the availability of clean, reliable, and comprehensive datasets is therefore essential for advancing research and practical applications.
+Another critical limitation is the lack of linguistic diversity.
+Existing datasets are available only in a few languages, which restricts their applicability to non-represented populations and limits the potential for cross-linguistic comparisons.
+This linguistic bias poses a challenge for developing universal models that can generalize across languages.
+In addition to these issues, most datasets are relatively small, often including only a limited number of participants.
+This scarcity hinders the development of robust and generalizable models and makes it difficult to capture the wide variability in speech disorders across different individuals and conditions.
+Furthermore, demographic imbalances are common.
+Many datasets exhibit skewed gender representation, which may introduce biases into trained models.
+In some cases, age groups or types of pathological conditions are also underrepresented, further limiting generalizability.
+These limitations underscore the pressing need for larger, more diverse, and better-balanced pathological speech datasets that reflect the complexity of real-world populations.
+Addressing these gaps is essential for building reliable and inclusive tools for pathological speech analysis.
+III Speech Representations for
+Automatic Approaches
+Pathological speech exhibits a range of acoustic anomalies, including deviations in pitch, loudness, vowel space reduction, and articulation [72, 4].
+Additionally, it can lead to asymmetrical tension in the vocal folds, resulting in irregular vibrations and, consequently, an abnormal fundamental frequency.
+Patients can also show inconsistent rhythmic structures in comparison to control groups [73].
+Excessively high or low fundamental frequency, combined with excessive vocal intensity, can exacerbate the severity of pathological voice conditions, producing characteristics such as shrillness, screechiness, hoarseness, or huskiness [74].
+To capture such abnormalities as biomarkers for automatic pathological speech processing, researchers have employed various handcrafted acoustic features such as OpenSMILE [70], Mel-frequency cepstral coefficients (MFCCs) [75], or spectro-temporal sparsity features [76].
+Raw speech signals and various time-frequency representations such as the short-time Fourier transform (STFT) have also been directly exploited in combination with deep learning approaches to directly learn pathology-discriminant cues [77, 78].
+More recently, latent embeddings derived from self-supervised models have been used as more powerful representations of speech patterns and the various impairments [70, 79, 80].
+It is important to note that a considerably larger set of features and representations have been employed to characterize disorders across various pathologies than the ones outlined above.
+Here, we briefly discuss the representations that are most commonly used in the literature.
+III-A Handcrafted Acoustic Features
+OpenSMILE
+Among the various features used in pathological speech detection, OpenSMILE features have been widely explored in the literature in conjunction with traditional machine learning algorithms [81, 82, 83, 84, 40, 85, 86, 87, 88, 89, 70, 90, 91, 92].
+The OpenSMILE feature set includes a $6552$-dimensional feature vector that primarily consists of low-level audio features such as CHROMA, CENS, loudness, MFCCs, and other spectral features.
+In the context of pathological speech, these features can capture the subtle anomalies in speech patterns that are indicative of disorders.
+OpenSMILE’s comprehensive feature set allows for the analysis of prosodic elements like pitch, jitter, shimmer, and formant frequencies, which are crucial for identifying and differentiating various speech pathologies.
+However, OpenSMILE features are general features that have been used for a variety of speech applications such as emotion recognition [93] or speech recognition [94] and they are not specifically handcrafted to capture pathological cues..
+Spectral and cepstral coefficients
+Due to their ability to characterize articulation deficiencies, various spectral and cepstral coefficients such as Linear Predictive Coding (LPC), MFCCs, and Perceptual Linear Prediction (PLP) features, have been successfully used for pathological speech detection [95, 96, 97, 98, 99, 100, 101, 102, 103, 13, 104].
+LPC features describe the distribution of energy across frequency bands and are directly related to the resonant properties of the vocal tract, making them suitable for analyzing articulatory features like formant frequencies and tongue positioning.
+In contrast, MFCCs and PLP features are obtained by transforming the spectral envelope into the cepstral domain, which effectively separates source and filter components while capturing smoothed representations of the vocal tract shape.
+When extended with temporal dynamics, these features can track transitions and instability in articulation patterns.
+To enhance robustness against channel variability and noise, the Relative Spectral Transform PLP features have also been considered in [105, 106].
+Furthermore, the fusion of modulation spectra features or glottal features with MFCCs has additionally been explored to further improve performance [107, 108].
+Spectro-temporal sparsity
+Since pathological speech can be breathy, semi-whispery, and is characterized by abnormal pauses and imprecise articulation, it can be expected that its spectro-temporal sparsity differs from the spectro-temporal sparsity of neurotypical speech.
+To characterize spectro-temporal sparsity, various sparsity-based features have been introduced in [76, 109].
+Although such features have been shown to be discriminative of various speech disorders[76, 109, 92, 86], they are highly sensitive to environmental artefacts such as noise and reverberation.
+III-B Time-Frequency Representations
+Input representations such as the STFT and its variants allow for the analysis of speech signals in both time and frequency domains, providing valuable insights into speech characteristics such as pitch, formant shifts, and spectral irregularities.
+The STFT is often employed due to its ability to capture dynamic changes in the signal over time, which is critical for identifying variations in speech patterns linked to speech disorders [110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 34, 120, 121].
+Furthermore, wavelet transforms and the Continuous Wavelet Transform offer better time-frequency localization, which is especially beneficial for analyzing transient and non-stationary features of pathological speech [122].
+III-C Raw Waveform Representation
+While handcrafted acoustic features and time-frequency representations have demonstrated strong performance, researchers have also explored end-to-end pathological speech detection using raw input representations, eliminating the need for handcrafted features or time-frequency representations.
+Studies such as [78, 123, 124] have shown that raw waveform-based methods can capture discriminative pathological patterns and outperform traditional feature-based approaches.
+However, these methods may require more training data compared to other approaches for a robust performance.
+III-D Self Supervised Embeddings
+Even though handcrafted acoustic features, time-frequency, and raw waveform input representations have achieved promising results in the analysis of pathological speech, their performance remains limited.
+The current state-of-the-art input representations for pathological speech primarily consist of latent embeddings derived from self-supervised models such as wav2vec2 [125], HuBERT [126], or WaveLM [127].
+Research in e.g., [128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 79, 141, 142, 143, 144, 92, 145, 146, 147, 148, 149, 150, 151] has shown that these advanced models are effective in capturing complex speech patterns and nuances, leading to improved performance in tasks like detection and recognition.
+However, ongoing research continues to refine these embeddings and assess their robustness and generalizability across diverse pathological speech conditions, tasks, and applications.
+IV Automatic Pathological Speech Detection
+As discussed in Section I, clinicians traditionally rely on laborious and time-consuming auditory-perceptual measures to diagnose speech impairments accurately.
 To address the various challenges associated with auditory-perceptual assessments, there has been a growing interest in the research community to develop automatic approaches for diagnosing pathological speech.
-The goal of these approaches is to enhance the detection accuracy to match or surpass human accuracy, thereby ensuring more consistent and reliable diagnoses.
-Automated systems leverage advanced technologies such as classical machine learning and deep learning models with input representations as discussed in Section II to analyze speech patterns and identify anomalies indicative of these neurodegenerative conditions.
+The goal of these approaches is to enhance the detection accuracy to match or surpass human accuracy, thereby ensuring more consistent and reliable diagnosis.
+Automated systems analyze speech patterns and identify anomalies indicative of these neurodegenerative conditions by leveraging classical machine learning and deep learning models with input representations such as the ones discussed in Section III.
 In the following, we briefly summarize various classical machine learning-based and deep learning-based approaches.
-Additionally, insights on the speech modes used for pathological speech detection are provided.
 IV-A Classical Machine Learning-based Approaches
-[126, 127] employed random forests and support vector machines (SVM), reporting promising results on a small pool of speakers.
-Similarly, [128] demonstrated good performance in detecting pathological dysarthric speech in an Indian Tamil language dataset using Gaussian mixture model and hidden Markov model classifiers.
-[129] found that the use of articulatory features on a small set of 24 Czech speakers showed very good performance, but might not generalize well to larger sets of speakers.
-Despite the promising results from these methods using various spectral/cepstral features, it remains unclear which features are most discriminative in capturing pathological cues.
-[130] showed that the cepstral coefficients are more robust and achieved the highest performance on a Spanish dataset when used as input to the SVMs for detection of pathological speech.
-[131] exploits $63736373$ acoustic, $36003600$ articulatory, and $44$ sensory features for the detection of ALS using deep neural networks and showed state-of-the-art detection performance on a dataset consisting of $1111$ patients and $1111$ healthy speakers.
-In a similar fashion, [132] employed SVM with openSMILE features on a dataset of 123 Hebrew speakers.
-Additionally, [133] utilized glottal features and demonstrated that combining glottal features with openSMILE features enhances the performance of dysarthric speech detection.
-[45] introduced spectro-temporal sparsity features combined with SVMs, showing these features to be more effective discriminators for dysarthric speech compared to temporal sparsity features.
-[134] applied Grassmann discriminant analysis on spectro-temporal subspaces, using singular value decomposition based on clinical evidence of spectro-temporal distortions.
-Their models were evaluated on datasets-PC-GITA and MoSpeeDi.
-The major limitation of such works is their focus on small speaker sets and bias towards factors such as gender, recording conditions, age, class, or language, leading to poor generalization in clinical scenarios.
-To investigate the impact of the distribution shift, [135] employed cross-database models for dysarthric pathological speech detection.
-They demonstrated that an SVM model trained on the UA-Speech dataset showed a significant performance drop when tested on the AMSDC dataset.
-This highlights the need for future work to develop models robust enough to handle distribution shifts, including variations in recording conditions and other factors.
-Furthermore, classical machine learning-based methods primarily rely on hand-crafted features, which may not effectively capture all abstract pathological cues.
-Finally, these methods also tend to overlook valuable metadata such as speaker identity, gender, and language.
+In recent years, numerous studies have investigated the use of classical machine learning classifiers combined with hand-crafted acoustic features for detecting pathological speech, yielding promising results.
+For instance, [152] and [153] employed Random Forests and Support Vector Machines (SVMs) on a range of dysphonia measures, achieving strong classification performance on relatively small speaker datasets.
+Similarly, [154] applied Gaussian Mixture Models and Hidden Markov Models to cepstral features for detecting dysarthric speech in an Indian Tamil language dataset.
+Orozco‐Arroyave et al.
+[155] further demonstrated that cepstral coefficients are effective and robust when used with SVMs for detecting dysarthric speech in a Spanish dataset.
+While these findings are encouraging, their scalability to larger, more diverse datasets remains uncertain due to the variability of speech pathologies across populations.
+Additionally, despite the effectiveness of various spectral and cepstral features, the field still lacks consensus on which features are the most discriminative and generalizable across different conditions and datasets.
+Recognizing the importance of developing discriminative feature representations, the research community has devoted significant efforts to this area.
+For example, [156] investigated the use of articulatory features for pathological speech detection and achieved promising results on a small dataset of 24 Czech speakers.
+However, the limited sample size restricts the generalizability of the findings, and extracting articulatory features at scale remains technically challenging.
+In a more comprehensive effort, [157] utilized an extensive feature set including $6,373$ acoustic, $3,600$ articulatory, and $4$ sensory features to detect ALS, reporting state-of-the-art performance on a dataset of $22$ speakers.
+Norel et al.
+[158] employed SVMs with openSMILE features to detect pathological speech in a larger cohort of 123 Hebrew speakers.
+Prabhakera and Alku [82] found that incorporating glottal features alongside openSMILE features improved dysarthric speech detection performance.
+Further advancements were made in [76] by introducing spectro-temporal sparsity features, which outperformed temporal sparsity features in classifying dysarthric speech using SVMs.
+Additionally, [159] applied Grassmann discriminant analysis to spectro-temporal subspaces, leveraging singular value decomposition informed by clinical insights into pathological distortions.
+A common limitation across these studies is their reliance on relatively small speaker sets and their susceptibility to biases related to sex, age, recording conditions, or language.
+To evaluate generalization across different domains, [160] conducted a cross-database study, demonstrating that an SVM trained on one dataset suffered a significant performance drop when evaluated on another dataset.
+This highlights the critical need for models that are robust to distributional shifts, including those stemming from differences in dataset characteristics and recording environments.
+Moreover, classical machine learning approaches heavily depend on hand-crafted features, which may not fully capture the nuanced and abstract cues associated with pathological speech.
+These methods also tend to overlook important metadata such as speaker identity, sex, and language.
 IV-B Deep Learning-based Approaches
-[136, 30, 137] utilized convolutional neural networks (CNN) for dysarthric speech detection on a dataset consisting of 26 speakers, including 13 patients diagnosed with ALS and their gender-matched counterparts respectively.
-Although this method demonstrated promising results, there is significant variation in pathological detection at the individual speaker level.
-[46] employed phase spectra such as modified group delay and instantaneous frequency spectra, demonstrating that using these phase spectra as complementary features to magnitude spectra improves the performance of CNN models in detecting dysarthric speech.
-Similarly, [56] employed temporal envelope features with CNN for dysarthric classification.
-[138] used SVM and deep neural networks for the detection of pathological speech on a dataset of speakers diagnosed with ALS.
-While most proposed approaches are single-task approaches aiming at achieving only pathological speech detection, [139] introduced the first ever multi-task system where they trained jointly pathological classification (primary task) and other 11 articulatory deficit tasks.
-The representations learned during joint training are constrained by various speech tasks and shows better generalization in pathological speech detection across speakers on a Spanish dataset.
-Most studies to date have not exploited the temporal nature of audio frames.
-[140, 51, 141] addressed this issue by introducing a long short-term memory (LSTM) temporal network for dysarthric speech classification.
-[140] evaluated LSTM on a Mandarin dataset comprising of 69 speakers, 31 diagnosed with dysarthria and the remaining as normal speakers, while [51] evaluated an LSTM-based model on the TORGO dataset.
-Although the methodology in [140] showed promising results, the model has been evaluated only on the syllables making it harder to generalize across spontaneous speech.
-Despite the promising results achieved by the proposed methods using handcrafted acoustic features and time-frequency representations, their detection capabilities still fall short of human-level performance and remain inadequate for clinical applications.
-Instead of handcrafted acoustic features or time-frequency representations, researchers endeavored to leverage raw speech data and applied self-supervised learning models like w2v2 to derive meaningful contextual representations, with the goal of improving the detection performance of pathological speech [60, 142, 43].
-To the best of our knowledge, such models have become the state-of-the-art models in pathological speech detection, outperforming other automatic approaches based on handcrafted acoustic features or time-frequency input representations.
-However, although advantageous, these models currently lack interpretability insights that might be useful for a clinician.
-Moreover, the models developed thus far are largely language-dependent, as the linguistic context plays a critical role in pathological speech.
-Developing language-agnostic pathological detection methods would be highly beneficial, allowing for universal application across patients worldwide [143].
-IV-C Impact of Speech Mode on Pathological Speech Detection
-The large majority of the previously reviewed automatic pathological speech detection approaches have been proposed and validated on controlled speech tasks.
-Controlled speech, also known as non-spontaneous speech, involves utterances produced within a structured context.
-In this mode, neurotypical speakers and patients are given a phonetically balanced, carefully crafted text and asked to repeat it.
-The goal is to elicit pathological biomarkers and cues through these controlled utterances, which can offer insights into motor planning and execution.
-Tasks in this mode may include reading from a text and repeating specific words.
-In contrast to controlled speech, non-controlled or spontaneous speech consists of unplanned utterances that mirror real-world communication.
-This type of speech requires greater cognitive effort for planning, motor execution, and articulation, making it more likely to reveal natural pathological biomarkers.
-Examples include storytelling, recounting incidents, or engaging in everyday conversations.
-The reliance on controlled speech, with its structured and phonetically balanced content, may not always capture the full spectrum of pathological cues.
-Incorporating spontaneous speech could provide a more comprehensive understanding of pathological speech patterns.
-It seems plausible that developing pathological speech detection methods for spontaneous speech could be highly beneficial, given its ease of collection [61].
-Research in this area could make such methods more practical and effective, as spontaneous speech naturally reflects real-world communication and may offer a richer set of pathological cues.
+Several studies have explored CNN-based architectures for pathological speech detection, leveraging their ability to extract local patterns from time-frequency representations.
+For instance, [39, 161] applied CNNs to dysarthric speech, achieving promising results on small datasets of control speakers and ALS patients.
+However, these models exhibited significant inter-speaker variability in performance, suggesting a lack of robustness across individuals.
+This limitation is particularly critical in clinical contexts where model reliability must generalize across diverse patient populations.
+Building on CNN-based models, [77] incorporated phase-based features, i.e., the modified group delay and instantaneous frequency spectra, as complementary inputs to magnitude spectra.
+This multi-representational approach improved the CNN performance for pathological speech detection.
+Similarly, [88] introduced temporal envelope features, reinforcing that temporal dynamics encode critical pathological cues.
+However, these representation augmentations increase model complexity and demand careful preprocessing, which may hinder real-world deployment.
+While most approaches focus solely on binary classification of pathological versus neurotypical speech, multi-task learning has shown potential for capturing broader articulatory patterns.
+Vásquez Correa et al.
+[162] introduced a multi-task system based on CNNs that simultaneously learned to classify pathological speech and predict $11$ articulatory deficit attributes.
+This approach yielded improved generalization across speakers by constraining the learned representations with related speech tasks, demonstrating the utility of auxiliary objectives in enhancing model robustness.
+While CNNs excel at spatial pattern recognition, they lack mechanisms to capture temporal dependencies inherent in speech.
+Recurrent architectures, particularly LSTMs, have been employed to address this gap [163, 83, 164, 123].
+These models demonstrate improved performance but often require larger datasets to train effectively.
+Moreover, although the LSTM methodology in [163] showed promising results, the model has been evaluated only on syllables and its generalization to other speech modes remains unexplored.
+More recently, there has been a shift toward end-to-end models using self-supervised learning (SSL) to bypass manual feature engineering.
+SSL models like wav2vec 2.0 leverage raw waveform inputs to learn contextual embeddings, achieving state-of-the-art results across several pathological speech datasets [91, 165, 70].
+These models outperform classical approaches in both accuracy and scalability.
+However, despite their empirical success, SSL-based models function as black boxes, raising concerns about clinical interpretability and decision transparency.
 V Pathological Speech Recognition
-Automatic speech recognition (ASR) systems have significantly advanced over the years, demonstrating impressive results in converting raw speech signals into their corresponding textual form.
+ASR systems have significantly advanced over the years, demonstrating impressive results in converting raw speech signals into their corresponding textual form.
 This has resulted in the widespread use of various interactive devices, including smartphones and voice assistants.
 However, they often fail to recognize low resource pathological speech, including speech from patients suffering from various neurodegenerative conditions such as PD or ALS [31].
-To alleviate this problem, several attempts have been made in improving the performance of ASR systems for low resource pathological speech.
+To alleviate this problem, several attempts have been made in improving the performance of ASR systems for pathological speakers.
 For example, [166] proposed a two-step speaker adaptation method.
-Initially, a model trained on extensive neurotypical speech data is fine-tuned for dysarthric ASR.
-Subsequently, this fine-tuned model undergoes further adaptation to the specific dysarthric speaker.
-Additionally, [145] proposed the additive angular margin loss to address intra-class variation among dysarthric speakers, demonstrating promising results with Japanese speakers.
+In the first step, a model trained on extensive control speech data is fine-tuned for dysarthric ASR.
+In the second step, this fine-tuned model undergoes further adaptation to a specific dysarthric speaker.
+Additionally, [167] proposed the additive angular margin loss to address intra-class variation among dysarthric speakers, demonstrating promising results on Japanese speakers.
+Green et al.
 [168] also showed that personalized ASR systems fine-tuned on pathological speech exhibit better recognition performance compared to speaker-independent ASR systems.
 However, the reliance on specific speakers in personalized ASR systems poses challenges for generalization settings.
-[147] utilize lattice-free maximum mutual information to mitigate insertion errors, which are otherwise prevalent due to the slow speaking rates of individuals with dysarthria.
-Due to the limited amount of available pathological speech data, the progress of ASR for pathological speech remains restricted.
-In a different approach, [148] utilized transfer learning and found that speaker-based data selection leads to negative transfer.
+Hermann and Magimai.-Doss [169] utilize lattice-free maximum mutual information to mitigate insertion errors, which are otherwise prevalent due to the slow speaking rates of individuals with dysarthria.
+In a different approach, Xiong et al.
+[170] utilized transfer learning and found that speaker-based data selection leads to negative transfer.
 They recommended using utterance-based data selection with an entropy distribution to enhance recognition.
+Yue et al.
 [171] further implemented a multi-task learning (MTL) framework through auto-encoder joint learning, utilizing bottleneck features on out-of-domain data.
 Employing MTL in pathological ASR showed a lower word error rate compared to its single-task counterpart.
-[150] demonstrated that state-of-the-art ASR systems for pathological speech are significantly impacted by phoneme inaccuracies.
-To address this, they proposed Speech Vision—a transfer learning paradigm that converts word utterances into visual feature representations, aiming to recognize the shape of the word rather than relying on phonemes.
-A comprehensive overview of ASR systems for pathological speakers in terms of progress and challenges is provided in [23].
-Due to the large number of parameters in state-of-the-art ASR models, fine-tuning them on low-resource pathological speech data is highly expensive.
-However, leveraging fine-tuning techniques such as LoRA (Low-Rank Adaptation) [151, 152] offers a more efficient alternative, reducing computational costs and making training more feasible.
-Additionally, LoRA’s modular approach might possibly make it easier to adapt the fine-tuned models to new speakers, providing flexibility in real-world applications or clinical scenarios.
+Shahamiri [172] demonstrated that state-of-the-art ASR systems for pathological speech are significantly impacted by phoneme inaccuracies.
+To address this, they proposed Speech Vision, a transfer learning paradigm that converts word utterances into visual feature representations, aiming to recognize the shape of the word rather than relying on phonemes.
+A comprehensive overview of ASR systems for pathological speakers in terms of progress and challenges is provided in [32], where it is shown that pathological speech recognition performance can be significantly improved through a combination of neural architecture search, data augmentation, speaker adaptation, and multi-modal learning.
+These techniques address challenges such as limited training data, high inter-speaker variability, and reduced speech intelligibility.
+It should be noted that due to the large number of parameters in state-of-the-art ASR models, fine-tuning them on low-resource pathological speech data can be highly expensive.
+However, leveraging fine-tuning techniques such as Low-Rank Adaptation (LoRA) offers a more efficient alternative, reducing computational costs and making training more feasible [173, 174].
+Additionally, LoRA’s modular approach might make it easier to adapt the fine-tuned models to new speakers, providing flexibility in real-world applications or clinical scenarios.
 VI Intelligibility Enhancement
 of Pathological Speech
-In pathological speech, enhancement refers to the process of improving the quality and intelligibility of speech samples affected by various impairments.
-The benefits of pathological speech enhancement are manifold.
-Firstly, the improved quality of pathological speech can be integrated into ASR systems, potentially enhancing the recognition accuracy of such speech, thereby increasing the usability of these systems for individuals with speech impairments.
-Secondly, these approaches can serve as data augmentation methods (r.f.
-VIII) for multiple applications such as pathological speech detection and automatic speech recognition.
-Progress in pathological speech enhancement is limited, however, several attempts have been made by various researchers as described in the following.
-[24] restructured the trajectory of formants to more closely align with the desired sound targets, thereby enhancing the accuracy of the acoustic representation.
-This refined formant trajectory was subsequently utilized to generate more intelligible speech, significantly improving the clarity and quality of the unintelligible pathological speech.
-[153] utilized a technique where the phase of the source signal was varied across six microphone arrays to synthesize pathological speech utterances.
-This approach effectively increased the data samples, which were later used for data augmentation in dysarthric ASR.
-[25] implemented a pause insertion model along with a severity level coefficient in a neural multi-task text-to-speech system, effectively replicating and synthesizing dysarthric speech.
-The synthesized speech was then utilized as data augmentation for improving dysarthric speech recognition.
-[154] enhanced the intelligibility of dysarthric speech by applying a methodology that includes pronunciation correction, insertion of missing sounds, devoicing of unvoiced phonemes, tempo and spectral adjustment, and removal of unwanted disfluencies.
-[155] demonstrated that modifying the short-term spectral content on a frame-by-frame basis can significantly improve the word-level intelligibility for patients with dysarthria.
-However, despite the substantial intelligibility improvements, the method’s speaker dependency poses challenges for generalizing across different speakers.
-In a similar speaker-dependent study, [156] employed a Kepstrum-based approach.
-[157] utilized multiple pronunciation patterns along with genetic algorithms and hidden Markov models [26, 27] to enhance dysarthric speech.
-[158] implemented a two-stage framework in which dysarthric speech was initially recognized using an hidden Markov-based network.
-Subsequently, a synthesis system adapted to the dysarthric speaker was employed to produce a new speech output by correcting the mispronounced words, thereby enhancing the intelligibility of the dysarthric speech.
-[159] used dynamic time warping with CNNs to improve the speech intelligibility.
-[160, 161] utilized linear predictive coding mapping in conjunction with dynamic time warping between dysarthric speech and neurotypical speech to enhance dysarthric speech intelligibility in an auditory feedback mechanism.
-[162] employed feature space mapping through constrained maximum likelihood linear regression, aligning acoustic features from dysarthric speakers to those of neurotypical speakers.
-This approach demonstrated improved intelligibility of dysarthric speech, as evidenced by the perceptual evaluation of speech quality test.
-A similar study was conducted by [76], where dysarthric speech features were enhanced using time-delay neural network denoising autoencoders in a speaker-independent fashion.
-However, only a small subset of speakers were used in the evaluation, which may not generalize well to a larger pool of speakers.
-Motivated by the human auditory perception system, [163] introduced perceptually enhanced single frequency cepstral coefficients to improve the intelligibility of speech in dysarthric patients using i-vectors and a Gaussian mixture model with a universal background model.
-[30] examined the spectro-temporal representation and showed it can be useful in improving the intelligibility of speech.
-[31] argued that prior methods for speech intelligibility enhancement primarily focus on speaker-dependent systems, and they introduced speech posterior feature sets designed to capture dysarthric attributes while ignoring speaker-specific characteristics.
-However, this assertion is not entirely accurate, as recent research has demonstrated that dysarthric speech is closely intertwined with speaker attributes [75].
-In a different approach, [164] integrated Transformer-TTS, CycleVAE-VC, and LPCNet to synthesize intelligible and robust dysarthric speech, although the resulting speech lacked naturalness.
-[28] formulated pathological speech enhancement as a style transfer problem and proposed a generative adversarial network (GAN)-based method for dysarthric speech enhancement.
-Similarly, [29] used CycleGAN for converting dysarthric speech to a more intelligible speech.
-[165] introduced DiscoGAN to improve the speech intelligibility of dysarthric patients.
-This method converts pathological speech to neurotypical speech through feature mapping learned during DiscoGAN training with an adversarial and construction losses [165].
-The approach showed promising results, though further validation with a larger and more diverse patient group is necessary to ensure generalizability across various pathological conditions.
-[166] explored a related approach and found that time-stretching techniques for dysarthric speech outperform existing state-of-the-art generative models.
-Their study concluded that using time-stretching with MaskCycleGAN achieves significantly better performance in dysarthric speech enhancement.
-However, they noted that despite these improvements, the method is not yet suitable for practical applications.
-Dysarthric speech datasets typically have a limited vocabulary, but using a conversion model can expand the set of out-of-vocabulary words, which can then be used for data augmentation.
-[167] presented the first end-to-end voice conversion system that trains dysarthric speech along with its corresponding textual form using a knowledge distillation approach.
-This method enables the replication of speech from dysarthric speakers with improved pronunciation and more precise articulation.
-To expand out-of-vocabulary word set, [78] utilized a voice conversion model to transform natural speech into synthetic dysarthric speech utterances.
-A detailed study of pathological voice conversion can be found in [168].
-Neural encoder-decoder models have been shown to significantly outperform GAN-based systems in enhancing the intelligibility of dysarthric speech.
-However, their reliance on multi-step modules often results in inefficient and suboptimal perceptual quality.
-To address this issue, [169] proposed Unit-DSR, a self-supervised learning approach for dysarthric speech synthesis.
-It has been shown that the adversarial training framework is highly sensitive to training parameters, often leading to instability and model divergence[170, 171].
-In contrast, diffusion models [172] show significant promise for applications in pathological speech enhancement, synthesis, and voice conversion.
-On a similar lines,  [173] employed the Glow-TTS diffusion model for generating dysarthric speech and used that as data augmentation for training a dysarthric ASR system.
+Pathological speech enhancement refers to improving the intelligibility and quality of speech affected by impairments such as dysarthria.
+The benefits of such enhancement are twofold.
+First, enhanced speech can ease human-human and human-machine communication for pathological speakers, promoting their social and digital inclusion.
+Second, enhancement methods can be leveraged for data augmentation (cf.
+Section VIII), aiding in the development of robust models for pathological speech processing.
+Early efforts in pathological speech enhancement focused on explicitly correcting articulatory and acoustic deficiencies.
+For instance, [33] improved intelligibility by restructuring formant trajectories to better match intended speech targets.
+Rudzicz [175] applied a range of techniques including pronunciation correction, phoneme insertion, tempo adjustment, and removal of disfluencies.
+Hosom et al.
+[176] modified short-term spectral features to enhance word-level intelligibility, though their approach was speaker-dependent.
+Lalitha et al.
+[177] proposed a speaker-specific Kepstrum-based method, while [178] introduced a two-stage framework that combined ASR with speech synthesis to correct mispronunciations.
+Several mapping-based approaches aiming to transform dysarthric speech into more intelligible forms have also been explored.
+For example, [179] proposed a speech enhancement method where a CNN is trained to directly map dysarthric utterances to their control counterparts.
+At the feature level, techniques such as LPC mapping and frequency warping of LPC poles have been explored in [180] and [181].
+Additionally, [111] investigated feature-level mapping using time-delay neural networks.
+More recent advances leverage voice conversion models for pathological intelligibility enhancement.
+For instance, [37] framed the task as a style transfer problem and employed GANs to convert dysarthric to typical speech.
+Similarly, [38] utilized CycleGAN for dysarthric-to-typical speech conversion, and [182] applied DiscoGAN to map pathological and typical speech features at the acoustic level.
+Among these, [183] showed that time-stretching combined with MaskCycleGAN outperformed other GAN-based models in intelligibility enhancement, although the technique’s practicality remains limited due to computational complexity.
+In a more comprehensive system, [184] integrated Transformer-TTS, CycleVAE-VC, and LPCNet to generate highly intelligible dysarthric speech, albeit at the expense of naturalness.
+Additionally, [185] proposed an end-to-end voice conversion system using knowledge distillation for enhanced dysarthric speech synthesis.
+To avoid adversarial training instability in GAN-based voice conversion models, neural encoder-decoder architectures have also been investigated for improving pathological speech intelligibility.
+For instance, [186] introduced Unit-DSR, which converts dysarthric speech into discrete linguistic units and then reconstructs speech from these normalized units using a neural vocoder.
+In summary, the evolution of pathological speech enhancement methods reflects a progression from early signal processing and mapping approaches to advanced generative models and neural architectures.
+While GAN-based models have significantly improved intelligibility, challenges remain in balancing enhancement quality, naturalness, and computational efficiency.
+Neural encoder-decoder frameworks show promise in addressing these challenges, marking an important direction for future research.
 VII Intelligibility and Severity Assessment
@@
 Non-blind approaches, by contrast, rely on intelligible speech from neurotypical speakers as a basis for estimating the intelligibility of pathological speech [194, 195, 196, 197, 198, 199, 200, 201, 202, 203].
-Such approaches typically use features extracted from ASR systems - which have been trained on large amounts of neurotypical speech - to train regression models to estimate the intelligibility of pathological speech.
+Such approaches typically use features extracted from ASR systems, which have been trained on large amounts of control speech, to train regression models to estimate the intelligibility of pathological speech.
 To avoid the burden of collecting and transcribing a large amount of neurotypical speech data required for such systems, [204] proposed the pathological short-time objective intelligibility measure (P-STOI) adapted from the speech enhancement domain.
-The P-STOI measure first calculates an utterance-dependent fully intelligible representation from a small set of neurotypical speakers.
+The P-STOI measure first calculates an utterance-dependent fully intelligible representation from a small set of control speakers.
 The intelligibility of the pathological utterance is then evaluated by quantifying its divergence from this reference representation in terms of the short-time spectral correlation.
-While advantageous, P-STOI requires recordings of the same utterance from intelligible neurotypical speakers, which may not always be available.
+While advantageous, P-STOI requires recordings of the same utterance from intelligible control speakers, which may not always be available.
 To mitigate this issue, [205] developed a method to generate synthetic reference speech for assessing pathological speech intelligibility.
-Likewise, [193] introduced a method for invariant reference speaker pair selection to evaluate speech intelligibility.
-This approach leverages the SpeechSplit framework, using dynamic time warping to measure the divergence between latent representations of healthy and pathological speech utterances.
 In a different approach, [206] introduced sub-space based intelligibility measures based on the premise that dominant spectral patterns in pathological speech deviate significantly from those of intelligible speech.
@@
 Besides intelligibility assessment, severity assessment is another important research area where developing tools for this purpose could greatly assist in automatizing the tedious process of screening patients and categorizing them into different subgroups based on the severity level.
-The methods developed in the literature for this purpose can be categorized into two categories, i.e., traditional machine learning-based approaches using the Mahalanobis distance classifier [174], SVMs [195], GMMs [196], or decision trees [197], and deep learning-based approaches [141, 198, 199, 200, 201, 59, 202, 203, 204].
-[205, 163] showed that i-vectors are effective in predicting the severity of pathological speech.
-Motivated by the human ear, [163] introduced perceptually enhanced single frequency cepstral coefficients for pathological speech intelligibility measure.
-Pathological speakers often exhibit irregular rhythm patterns in their speech.
-To leverage this characteristic, [206, 207] explored rhythm-based features for severity assessment, utilizing SVMs and MLP classifiers.
-Their study highlighted the potential of rhythm as a key indicator of pathological speech impairment severity, though challenges remain in capturing the full complexity of rhythm variations across different speakers and integrating these features into more advanced machine learning models.
-Furthermore, patients with various pathological conditions often display distinct facial expressions.
-To harness these visual features, [208] introduced the first audio-visual pathological severity classification system using convolutional neural networks.
-[200] introduced a novel training framework, demonstrating that incorporating weak speech supervision can offer significant advantages in pathological speech severity classification.
-To leverage metadata information such as age, gender, and type of pathological condition, [209] proposed a multi-head attention-based multi-task framework.
-This approach jointly optimizes severity, type, gender, and age classifications for pathological speech, thereby enhancing the robustness of latent features across these additional factors.
-Recently, there has been a shift towards leveraging self-supervised speech embeddings, such as Whisper and wav2vec2, in pathological speech severity measurement.
-This approach is gaining popularity due to the limited availability of pathological data and the ability to utilize unlabeled and metadata from other datasets, making it a promising direction for improving severity assessment in resource-constrained scenarios [77, 47].
+The methods developed in the literature for this purpose can be categorized into two categories, i.e., traditional machine learning-based approaches using the Mahalanobis distance classifier [187], SVMs [207], GMMs [208], or decision trees [209], and deep learning-based approaches [210, 211, 212, 90, 213, 214, 215].
+While deep learning approaches aim to automatically extract acoustic cues correlated with severity from raw or minimally processed speech signals, traditional machine learning approaches for severity assessment rely on (clinically informed) handcrafted acoustic features.
+For example, motivated by auditory processing knowledge, [216] introduced perceptually enhanced single frequency cepstral coefficients for assessing the severity of pathological speech.
+Vásquez-Correa et al.
+[217] showed that articulation features extracted from continuous speech signals to create i-vectors were advantageous in quantifing the dysarthria severity level.
+Based on the knowledge that pathological speakers often exhibit irregular rhythm patterns in their speech, [218, 219] explored rhythm-based features for severity assessment.
+Besides speech impairments, patients with various pathological conditions often display distinct facial expressions.
+To harness these visual features, [220] introduced the first audio-visual pathological severity classification system using CNNs.
+To leverage metadata information such as age, sex, and type of pathological condition, [221] proposed a multi-head attention-based MTL framework.
+This approach jointly optimizes severity, type, sex, and age classifications for pathological speech, thereby enhancing the robustness of latent features across these additional factors.
+Recently, there has been growing interest in using SSL embeddings to measure pathological speech severity.
+This approach is promising due to the scarcity of labeled pathological data and the ability to leverage unlabeled data and metadata from other datasets, making it well-suited for resource-constrained settings [112, 79].
 VIII Data Augmentation for
 Pathological Speech Applications
-To mitigate overfitting in deep learning models dealing with the low resource pathological speech data, data augmentation techniques have been applied for various tasks.
-Existing approaches to data augmentation in pathological speech can be broadly classified into the following categories: traditional augmentation, voice conversion, text-to-speech (TTS) synthesis, and generative models.
-Drawing inspiration from the temporal and speed differences between pathological and neurotypical speech, [210, 211, 84] proposed a data augmentation technique that involves adjusting the temporal and speed characteristics of neurotypical speech to simulate pathological speech.
+To mitigate overfitting in deep learning models dealing with the low resource pathological speech data, data augmentation techniques have been used for various tasks.
+Existing approaches to data augmentation in pathological speech are broadly based on traditional strategies (such as incorporating noise, reverberation, or multiple datasets), perturbation strategies, voice conversion, and text-to-speech (TTS) synthesis.
+Takashima et al.
+[222] employed a strategy that involved combining diverse pathological speech data from multiple languages to increase the number of data samples.
+Their results indicate that such a traditional data augmentation approach can be advantageous for pathological ASR.
+Based on the temporal and speed differences between pathological and control speech, [223, 224, 119] have explored vocal tract length perturbation, tempo perturbation, and speed perturbation as data augmentation approaches for pathological ASR.
 Similarly, [225] introduced a data augmentation technique that adjusts the phonetic-level tempo of healthy speech to resemble atypical speech, and vice versa.
 Their findings demonstrated that the former approach is more effective for pathological ASR.
-[213] employed a strategy that involved combining diverse pathological speech data from multiple languages to increase the number of data samples.
-Their results indicated that this approach offers certain advantages in pathological ASR.
-Additionally, they demonstrated that integrating data from various sources can improve model performance by providing a richer and more varied dataset for training.
-Generative models like GANs and variational autoencoders are also commonly employed to generate pathological speech samples from healthy speech [28, 173].
-These synthetic samples serve dual purposes: they can be used for speech enhancement, improving the clarity and quality of pathological speech, and for expanding the dataset, thereby enhancing the diversity of training data.
-By increasing the variety of pathological speech examples, these models help in mitigating overfitting and improving the performance of deep learning models in tasks such as pathological speech detection and pathological ASR [28, 173, 214, 215].
-Dysarthric speech datasets typically have a limited vocabulary, but using a conversion model can expand the set of out-of-vocabulary words, which can then be used for data augmentation.
-To this end, [78] utilized a voice conversion model to transform natural speech into synthetic pathological speech utterances.
-[83] proposed a method incorporating temporal and spectral masking combined with time warping within a contrastive learning framework to develop robust speech representations for pathological ASR.
-[216] employed a TTS framework to generate synthetic dysarthric speech samples as an augmentation for pathological ASR.
-Although this approach showed some advantages, it faced challenges in generalizing to unseen speakers.
-This difficulty arose from the limited variability in pathological speech data, which was insufficient to fully capture the diverse characteristics of pathological speech.
-[107] investigated different data augmentation techniques in pathological ASR and demonstrated that GAN-based perturbation methods are more effective than other augmentation approaches.
-However, a comprehensive investigation of data augmentation methods in pathological speech detection has not yet been conducted.
-It would be particularly important to explore various data augmentation techniques on clean datasets such as PC-GITA, rather than on the noisy datasets such as TORGO, which has been the case for the previously reviewed literature.
-IX Other Challenges and Research Directions in Pathological Speech Processing
-This section describes the various challenges and possible research directions for pathological speech analysis.
-IX-A Robustness in Pathological Speech Detection
-The methods developed so far have typically been designed and tested under specific environmental conditions and for a particular language.
+More recently, voice conversion and TTS systems are also commonly employed to generate pathological speech from healthy speech [37, 226, 227].
+These synthetic samples serve a dual purpose, i.e., they can be used for speech enhancement, improving the clarity and quality of pathological speech; and they can be used for expanding the dataset, thereby enhancing the diversity of training data.
+Increased data diversity is beneficial to avoid overfitting and improve the performance of deep learning models in tasks such as pathological speech detection and pathological ASR [37, 226, 228, 229].
+Since pathological speech datasets typically have a limited vocabulary, using a voice conversion or TTS model can also be used to expand the set of out-of-vocabulary words [113].
+Recently, [147] investigated different data augmentation strategies in pathological ASR, demonstrating that GAN-based conversion methods are more effective than perturbation-based augmentation approaches.
+However, a comprehensive investigation of the advantages of all data augmentation strategies in various pathological speech processing tasks using multiple datasets is still lacking.
+IX Challenges and Research Directions in Pathological Speech Processing
+In clinical settings, the integration of automated speech processing systems for pathological speech analysis is crucial for advancing the diagnosis, therapy, and monitoring of various disorders.
+While standardized clinical scales are invaluable in perceptually evaluating pathological speech, the incorporation of automated systems can offer additional benefits in terms of objectivity, efficiency, and real-time feedback.
+Tools such as VoxTester [230], for instance, provide clinicians with quantifiable speech metrics, including articulation precision and speech rate, which help monitor disease progression and assess the effectiveness of interventions.
+However, to fully realize the potential of these tools, further research is needed to ensure that they can seamlessly integrate into clinical workflows, offering clinicians user-friendly, reliable, and interpretable outputs.
+Addressing issues such as clinician training, data security, patient consent, and adaptability to various clinical settings is crucial for the real-world adoption of these systems.
+Moreover, creating systems that can operate longitudinally, i.e., tracking speech performance over time to capture subtle changes in speech function, would significantly enhance therapeutic outcomes.
+We believe that addressing the challenges and research directions outlined in the remainder of this section will be key to enabling the seamless integration of these technologies into clinical practice in the future.
+IX-A Impact of Speech Mode
+The large majority of previously reviewed deep learning-based pathological speech approaches have been proposed and validated on controlled speech tasks.
+Controlled speech, also known as non-spontaneous speech, involves utterances produced within a structured context, typically requiring participants to repeat phonetically balanced, carefully crafted texts.
+This mode is designed to elicit specific pathological biomarkers by standardizing the motor planning and execution demands.
+Tasks may include reading aloud or repeating scripted phrases.
+In contrast, spontaneous speech consists of unplanned utterances, such as storytelling or casual conversation, which reflect real-world communicative behavior.
+It places greater demands on cognitive planning, articulation, and natural language generation, and may therefore reveal more authentic or varied pathological cues.
+While this distinction has received some attention in the context of detection, like in [92], its implications extend across other subfields such as severity assessment, intelligibility prediction, enhancement, and recognition.
+For instance, severity assessment models trained only on controlled speech may generalize poorly to real-world settings.
+Similarly, enhancement models may be tuned to the acoustic patterns of controlled tasks, failing to capture the variability in spontaneous speech.
+Incorporating spontaneous speech more broadly into pathological speech research could enhance practical utility and performance across these subfields.
+Given its ease of collection and alignment with natural communication, spontaneous speech provides a more suitable and informative basis for training and evaluating models [92, 231, 232].
+Future work should consider systematically analyzing the effect of speech mode across multiple application areas to ensure generalizability and real-world effectiveness.
+IX-B Robustness in Pathological Speech Detection
+The methods developed so far for automatic pathological speech detection have typically been designed and tested under specific environmental conditions and for a particular language.
 Each dataset used for pathological speech detection has its own unique variabilities in terms of both inter-speaker and intra-speaker differences.
@@
 Language is another important factor; separate models have been independently developed for different languages.
-However, there is a need for a more robust model that is language-agnostic (to a possible degree given that pathological cues might be different in different languages), age-agnostic, accent-agnostic, and gender-agnostic to enhance overall effectiveness.
+However, there is a need for a more robust model that is language-agnostic (to a possible degree given that pathological cues might be different in different languages), age-agnostic, accent-agnostic, and sex-agnostic to enhance overall effectiveness [151].
 Additionally, datasets such as TORGO are contaminated with noise.
 Researchers have shown that models trained on these datasets tend to learn environmental factors rather than focusing on genuine pathological cues [70].
-To the best of our knowledge, the PC-GITA database is currently the only clean pathological dataset available under a research license.
-The development of pathological speech detection models robust to environmental distortions has been very limited, with only a few small studies addressing this area [217, 114, 113].
+The development of pathological speech detection models robust to environmental distortions has been very limited, with only a few small studies addressing this area [233, 234, 71, 235].
 For example, [234] employed a test-time adaptation method to fine-tune pre-trained models on a validation set augmented with the test noise extracted from the test utterance.
@@
 Besides robustness to environmental distortions, adversarial robustness of pathological speech detection models is another important topic and research direction.
-[115] explored the impact of acoustically imperceptible adversarial perturbations on deep learning-based pathological speech detection models.
+The impact of acoustically imperceptible adversarial perturbations on deep learning-based pathological speech detection models has been explored in [71].
 Results revealed a high vulnerability of such models to adversarial perturbations, with adversarial training ineffective in enhancing robustness.
-IX-B Graph Neural Networks for Pathological Speech Detection
-While many automatic pathological speech detection methods have demonstrated remarkable performance, the exploitation of common attributes such as pathological cues, speaker characteristics, age, and gender across different speakers remains limited.
-A promising direction would be to approach the pathological speech detection problem as a semi-supervised node graph classification task via graph neural networks (GNNs).
-This could involve constructing an inter-speaker graph from the utterances of various speakers, where the graph connectivity would help form speaker clusters based on the presence or absence of pathological cues.
-This structural information can then be leveraged through various graph message-passing mechanisms to improve detection accuracy and provide deeper insights into pathological speech patterns.
-Initial work in this direction has been done in [218], where a graph convolutional neural networks has been proposed for Parkinson’s disease detection method by framing the task as a node classification problem.
-In this approach, a graph is constructed using w2v2 speech embeddings across different speech segments from various patients as nodes, with the edges capturing the relationships between these segments.
-We believe that in this context, a promising research direction is to explore the incorporation of domain knowledge, such as establishing edges based on factors like gender, age, and severity scale of the patients.
-IX-C Privacy
+IX-C Improving the Performance of Automatic Pathological Speech Analysis
+While many automatic pathological speech detection methods have shown remarkable performance, the exploitation of common attributes such as pathological cues, speaker characteristics, age, and sex across different speakers remains limited.
+A promising direction is to approach the pathological speech detection problem as a semi-supervised node graph classification task using graph neural networks (GNNs), as demonstrated in [231].
+This approach could involve constructing an inter-speaker graph based on utterances from various speakers, where the graph’s connectivity would help form speaker clusters based on the presence or absence of pathological cues.
+In this context, domain knowledge can be easily integrated by establishing edges based on factors such as sex, age, and the severity scale of the patients.
+Additionally, the recent availability of longitudinal multimodal data [61] enables novel applications of GNNs in disease monitoring.
+By representing the patient history as a temporal graph where nodes capture multimodal features (e.g., vocal biomarkers, facial expressivity) and edges encode their dynamic interactions, GNNs can model disease progression through evolving graph topologies.
+An additional area that remains under-explored in current research, potentially due to the lack of large datasets, is the adoption of Bayesian frameworks and generative approaches for pathological speech processing.
+Bayesian frameworks are particularly relevant in healthcare, as they allow the model to estimate uncertainties in its predictions, which is essential for high-risk settings where incorrect predictions can have serious consequences.
+These models are ideal for situations where speech data may vary considerably due to speaker differences or environmental factors.
+By providing probabilistic models, Bayesian methods can account for such variability, making them highly suitable for clinical applications where individual differences are pronounced.
+IX-D Privacy
 In pathological speech-based applications such as detection models or ASR systems, privacy-preserving solutions are crucial for safeguarding sensitive patient data.
 Since these systems process speech that reveals sensitive medical information, ensuring data security and confidentiality is of paramount importance.
-By employing privacy-preserving techniques - such as federated learning, differential privacy, and encryption - providers can ensure that individuals’ speech data is anonymized and never exposed to unauthorized parties.
+By employing privacy-preserving techniques such as federated learning, differential privacy, and encryption, providers can ensure that individuals’ speech data is anonymized and never exposed to unauthorized parties.
 This not only fosters trust among patients but also complies with stringent data protection regulations, thereby mitigating the risk of breaches.
 Developing privacy-preserving pathological speech processing systems where one must balance two conflicting goals, i.e., increasing the utility of the models while preserving the privacy of the users, remains an important challenging research directions.
-IX-D Multimodal Pathological Speech Analysis
+IX-E Multimodal Pathological Speech Analysis
 Most existing systems primarily focus on leveraging speech cues for detecting pathological speech.
 However, complementary cues may also exist in visual forms, such as facial expressions or lip movements, which can provide additional insights.
-To effectively utilize these visual cues, multimodal frameworks like AV-HuBERT could prove advantageous, as they integrate both audio and visual information, potentially enhancing the accuracy and robustness of pathological speech detection models.
-Furthermore, combining other modalities, such as electroencephalogram (EEG) brain signals and textual data (transcripts), could offer a more complete characterization of disorders.
-These multimodal systems may help in capturing underlying neural, visual or linguistic patterns associated with speech pathology, improving detection and diagnosis across diverse patient groups.
-X Explainability and Interpretibility
+To effectively utilize these visual cues, multimodal self-supervised methods such as AV-HuBERT [237] could prove advantageous, as they facilitate the integration of multimodal audio and visual information, potentially enhancing the accuracy and robustness of pathological speech detection models.
+Furthermore, combining other modalities, such as medical imaging, health records, brain signals, and textual data, could offer a more complete characterization of disorders.
+These multimodal systems may help in capturing underlying neural, visual or linguistic patterns associated with speech pathology, improving pathological speech processing across diverse patient groups.
+IX-F Explainability and Interpretability
 In the domain of pathological speech detection, explainability and interpretability are crucial due to their clinical significance.
-Although these terms are often used interchangeably because of their close overlap, it’s important to distinguish between them.
-As defined by [219], interpretable models are designed to be inherently understandable, whereas explainable models provide post-hoc explanations for the decisions made by existing ”black box” systems that are otherwise incomprehensible to humans.
-[220] further clarified the distinction by defining interpretability as ”How does the model work” and explainability as ”What else can the model tell me.” Keeping these distinctions in mind, particularly in the context of pathological speech detection, there has been very limited, if any, attention given to this aspect of research [221, 222, 223, 224].
-Most existing studies focus on improving model accuracy and performance, often overlooking the importance of making these models interpretable and explainable.
-This gap is critical, as the ability to understand and explain the decision-making process in pathological speech detection models can provide valuable insights for clinicians and support more trustworthy and reliable use of artificial intelligence in healthcare applications.
-To address this issue, [221] attempted to map high-dimensional acoustic features to sound pattern of English binary phonological features, recognizing that these high-dimensional acoustic features are challenging to interpret in clinical practice.
-The problem with this approach is that it may be biased toward English sounds, making it susceptible to errors when applied to different phonetic environments.
-Further, [224] applied the SHAP algorithm to identify the most influential features in their model and discovered that constant-vowel transitions were the key features in reversing the model’s decisions in pathological speech detection.
-Recently, [225] demonstrated using canonical correlation analysis that the human frequency range (0–210 Hz) plays a critical role in influencing model decisions for pathological speech detection.
-Although some research has been conducted, substantially more research is needed to address this under-explored but very important area.
-XI Large Language Models
-for Pathological Speech
-Given the recent advancements in large language models (LLMs), which have demonstrated significant progress across a wide range of applications [226, 227], exploring their potential for pathological speech analysis appears to be an inevitable and promising direction.
-LLMs could be leveraged for tasks such as detecting speech impairments based on specific phrase cues, improving the automatic classification of speech disorders, and enhancing the reconstruction of unintelligible or difficult speech.
-Additionally, LLMs’ ability to capture complex linguistic patterns may lead to more accurate models for personalized therapy, rehabilitation, and assistive communication tools for individuals with speech impairments.
+Although these terms are often used interchangeably, it is important to distinguish between them.
+As defined in [238], interpretable models are designed to be inherently understandable, whereas explainable models provide post-hoc explanations for the decisions made by existing black box systems that are otherwise incomprehensible to humans.
+Despite their importance, relatively little attention has been paid to explainability and interpretability in pathological speech detection, with most existing research focusing primarily on improving model accuracy.
+Nonetheless, a few notable efforts have begun to address this gap [239, 240, 241, 242, 243].
+For example, [239] mapped high-dimensional acoustic features to binary phonological representations, recognizing that raw acoustic features are difficult to interpret in clinical practice.
+Xu et al.
+[243] applied the SHAP algorithm to identify the most influential features in their model, highlighting the role of consonant-vowel transitions in reversing classification decisions.
+Similarly, [232] used canonical correlation analysis to show that the 0210210 Hz frequency range strongly influences model outputs.
+Recently, [151] found that non-interpretable SSL embeddings outperform interpretable features (e.g., prosodic, linguistic, and cognitive descriptors) in both multilingual and cross-lingual contexts.
+This highlights a growing challenge, i.e., as models increasingly rely on high-performing but opaque representations like SSL embeddings, there is a pressing need to develop methods that make these models explainable and clinically trustworthy.
+IX-G Large Language Models for Pathological Speech
+Given the recent advancements in multimodal large language models (LLMs), which have demonstrated significant progress across a wide range of applications [244, 245], exploring their potential for pathological speech analysis appears to be an inevitable and promising direction.
+Multimodal LLMs could be leveraged for tasks such as detecting speech impairments, ASR, and enhancing the reconstruction of unintelligible or difficult speech.
+Their ability to capture complex patterns may lead to more accurate models for personalized therapy, rehabilitation, and assistive communication tools for individuals with speech impairments.
+Additionally, multimodal LLMs should be explored for their ability to explain their decisions.
 This exploration could bridge the gap between traditional speech processing techniques and state-of-the-art language models, opening up new avenues for more effective and adaptive speech rehabilitation systems.
-XII Conclusion
-This paper provides a comprehensive overview of deep learning applications for pathological speech, encompassing detection, recognition, intelligibility assessment and enhancement.
-Additionally, it compiles a thorough list of both accessible and non accessible pathological speech datasets, which will serve as valuable resources for future research and accelerate progress in the field.
-To the best of our knowledge, this is the first review to offer an in-depth survey of pathological speech from the perspectives of detection, recognition, enhancement, explainibility and data augmentation.
-Furthermore, it outlines potential future research directions, particularly in the context of robust and interpretable models that can be deployed in clinical practice.
+X Conclusion
+This paper provides a comprehensive overview of speech analysis and technologies for pathological speech arising due to neurological disorders, encompassing detection, recognition, intelligibility assessment, and enhancement.
+Additionally, it compiles a thorough list of both accessible and non-accessible pathological speech datasets, which will serve as valuable resources for future research and accelerate progress in the field.
+Finally, it outlines potential future research directions, particularly in the context of robust and interpretable models that can be deployed in clinical practice.
 Acknowledgment